### v1

# Canonicalization for Unreproducible Builds in Java

Abstract.
The increasing complexity of software supply chains and the rise of supply chain attacks have elevated concerns around software integrity. Users and stakeholders face significant challenges in validating that a given software artifact corresponds to its declared source. Reproducible Builds address this challenge by ensuring that independently performed builds from identical source code produce identical binaries. However, achieving reproducibility at scale remains difficult, especially in Java, due to a range of non-deterministic factors and caveats in the build process. In this work, we focus on reproducibility in Java-based software, archetypal of enterprise applications. We introduce a conceptual framework for reproducible builds, we analyze a large dataset from Reproducible Central, and we develop a novel taxonomy of six root causes of unreproducibility. We study actionable mitigations: artifact and bytecode canonicalization using OSS-Rebuild and jNorm respectively. Finally, we present Chains-Rebuild, a tool that raises reproducibility success from 9.48% to 26.89% on 12,283 unreproducible artifacts. To sum up, our contributions are the first large-scale taxonomy of build unreproducibility causes in Java, a publicly available dataset of unreproducible builds, and Chains-Rebuild, a canonicalization tool for mitigating unreproducible builds in Java.
Reproducible Builds, Software Supply Chain, Canonicalization, Java
Figure 1. The problem space of unreproducible builds in Java, and our novel mitigation strategies.
1. Introduction
The growing complexity of software supply chains (Soto-Valero et al., 2022; Cox, 2019), coupled with the increasing frequency of supply chain attacks  [^1] https://www.sonatype.com/hubfs/SSCR-2024/SSCR_2024-FINAL-10-10-24.pdf, raises concerns about software integrity. In such a fragmented ecosystem, relying solely on assumptions about the identity of the distributor (Thompson, 1984) is no longer sufficient - what is needed is verifiable evidence that the software one installs corresponds exactly to its declared source. This challenge is especially acute in open source environments, where binaries are often distributed separately from their source code (Davies et al., 2011), making it difficult for users to independently validate what they are running. As a result, the focus is shifting toward techniques that can guarantee that what is built matches what was intended, regardless of who performs the build.
This technique is formally called “Reproducible Build” (Williams et al., 2025; Cox, 2025): builds are considered “reproducible” if and only if the build process is deterministic, where the same binary can be computed from the same source code by independent parties (Lamb and Zacchiroli, 2022). It helps prevent attacks on the build process, where an attacker modifies it to insert backdoors or malicious code into the software application to be built (Perry, 2013). Any backdoor or malicious code would be detected by the verifier as the built artifact would not match the original artifact. This approach has already seen adoption in security-critical environments, such as in certain federal government contracts, where vendors are required to supply source code so the agencies can perform their own builds and verify the integrity of the software  [^2] https://news.ycombinator.com/item?id=43492115.
Despite the promises of Reproducible Builds, ensuring and verifying reproducibility at scale remains technically challenging due to the multitude of spurious differences in build outputs (e.g., timestamps, file ordering). Those differences make binaries appear different even when built from the same source, with an untampered build pipeline (Fourné et al., 2023). Spurious differences hinder reproducibility and make it difficult to justify the differences between two builds. XZ Utils is a widely used compression library that is used in almost all Linux distributions. In 2024, it is found that the XZ Utils has a backdoor to give remote code execution to the attacker (Przymus and Durieux, 2025). However, this backdoor is only observed in the official tarballs on the registry and do not exist in the git repository. Such supply chains attacks can be prevented if build process is reproducible. In the case of XZ Utils, if the tarball on the package registry does not match the one built from the source code, then the user can be sure that the tarball is tampered with (Malka, 2025).
In this work, we contribute to solving the problem of achieving build reproducibility in the context of enterprise software in Java. Recall that Java has consistently been one of the top 5 programming languages in the world for the past 5 years  [^3] https://innovationgraph.github.com/global-metrics/programming-languages, underscoring its widespread use and relevance. Its importance in the context of finance (Soto-Valero et al., 2022; Morales, 2021), government services (Robinson, 2017), and military applications  [^4] http://pdf.cloud.opensystemsmedia.com/vita-technologies.com/Aonix.Jun07.pdf  [^5] https://tsri.com/news-blog/press/u-s-department-of-defense-mainframe-cobol-to-java-on-aws highlights the need for reproducible builds in Java. We aim at building the most comprehensive taxonomy of unreproducible builds in Java, and the corresponding mitigation strategies. Furthermore, we perform a deep study how canonicalization - removal of non-deterministic differences - is a good solution for mitigating unreproducibility.
Our approach for analyzing unreproducible builds in Java is shown in Figure 1. We first propose an original framework for build reproducibility where we clearly define the roles of the builder and rebuilder and how both of them contribute to reproducible build verification. Next, we build a dataset of unreproducible builds in Java by leveraging Reproducible Central (Boutemy, 2024), the leading project for rebuilding and verifying Maven applications. We systematically analyze the dataset to identify and classify the causes of unreproducibility into an original taxonomy of unreproducibility. Finally, we leverage the dataset to evaluate the effectiveness of bytecode and artifact canonicalization. Bytecode canonicalization focuses on internal representation of program logic and eliminates compiler introduced variations. While artifact canonicalization eliminates spurious variations in metadata that can influence how the artifact is interpreted or processed by downstream build tools. Together, these two levels of canonicalization address both semantic and non-semantic sources of non-determinism, significantly improving overall build reproducibility.
Our study reveals 6 original categories of unreproducibility causes in Java, each with the underlying root causes and the best practice mitigations. For example, changes in the bytecode in Java artifacts. Our large-scale experiments empirically measure the effectiveness of two state-of-the-art canonicalization techniques. First, we show that bytecode canonicalization with jNorm (Schott et al., 2024) can successfully canonicalize 267 out of 898 unreproducible artifacts with JVM bytecode. We study the reasons why jNorm fails to canonicalize the remaining artifacts, which are promising areas of future work. Next, we evaluate artifact canonicalization using Google’s OSS-Rebuild and find that it can successfully canonicalize 9.48% (1,165 / 12,283) of unreproducible artifacts. By implementing new canonicalization patterns informed by our taxonomy, our new tool Chains-Rebuild increases the verification success rate to 26.89% (3,303 / 12,283). Our paper is the first to demonstrate, at a large scale, that canonicalization is a promising mitigation strategy for fixing build unreproducibility in the Java ecosystem.
There has been prior work on reproducible builds in Java focusing on bytecode canonicalization (Dietrich et al., 2024b) and fixing build process (Xiong et al., 2022). However, we are the first to analyze Maven projects hosted on Maven Central at scale and to propose a comprehensive taxonomy of unreproducibility causes and their mitigations. In addition to bytecode canonicalization, we also study artifact canonicalization and show that it is a promising solution to mitigate unreproducibility in Java.
To summarize, our contributions are:
  • A sound conceptual framework of build reproducibility, founded on the novel roles of the “builder” and “rebuilder” and on the powerful concept of artifact canonicalization.
  • A novel, comprehensive taxonomy of 6 unreproducibility causes and their mitigation in Java, helping practitioners to improve build reproducibility in their projects.
  • A large-scale experiment on the effectiveness of canonicalization, showing that it is a promising solution to mitigate unreproducibility in Java. Artifact canonicalization is more effective at scale than bytecode canonicalization.
  • A publicly available, curated dataset of 7,961 unreproducible Maven releases for future research.
2. Framework for Build Reproducibility
2.1. Definitions
Developers often release reusable software in package registries, making them widely available for integration into third-party software projects. For our work on build reproducibility, we need to introduce key terminology. We work in the context of the Java programming language and its Maven build system.
Package Registry: A package registry is a centralized repository of software packages that can be reused by other applications. In this paper Maven Central, hosted by a company called Sonatype, is the considered package registry.
Module: A module is a uniquely addressable component in a software system. In Maven, a module is referenced by a combination of Group Identifier, Artifact Identifier, and Version. Group ID is a unique identifier for the organization, eg. “org.apache”   [^6] https://maven.apache.org/guides/mini/guide-reproducible-builds.html. Artifact ID is the unique name of the module and is unique within the group ID. For example, com.google.guava:guava:32.0.0-jre is a Maven module where com.google.guava is the group ID, guava is the artifact ID, and 32.0.0-jre is the version. There are two types of modules: a source module is a module that is hosted on a version control system (eg. Github), not necessarily published on a package registry; a released module is one published on a package registry, Maven Central in this paper. Both released and source modules can be identified by the combination of group ID, artifact ID, and version to refer to both types of module. In the latter case, we assume the presence of a mapping from version to Git tag or Git commit.
Project: A Maven project is a collection of source modules in the same source repository. It contains build configuration for all the modules in a so-called ‘POM file’. For example, Maven project guava contains 6 source modules - guava-parent, guava-tests, guava-testlib, guava-gwt, guava-bom, and guava.
Artifact: An artifact is any file that is part of the released module. In Maven, there are mostly a JAR or a ZIP archive. For example, in released com.google.guava:guava:32.0.0-jre, guava-sources.jar and guava.jar are JAR file in the released module. In this paper, we have two types of artifacts: The reference artifact is the artifact of the released module that is available on the package registry; The rebuild artifact is the artifact obtained by rebuilding the source module.
Builder: The builder is the entity that publishes a released module on a package registry, from the source module. A builder is often an automated script implemented in a CI/CD pipeline, like GitHub Actions. For example, the released module com.google.guava:guava:32.0.0-jre has been deployed on Maven Central by the Google builder.
Rebuilder: The rebuilder is the entity that verifies the build reproducibility of the released module, with information available from the package registry. In other words, the rebuilder 1) builds the projects again 2) verifies the reproducibility of all the rebuild artifacts against the artifacts in the released module. For example, the Reproducible Central project (Boutemy, 2024) is a rebuilder that verifies the reproducibility of the released modules of 700+ open-source Java libraries.
Canonicalization Canonicalization is the process of converting different representations of the same data into a single, standard representation. For example, in the domain of XML parsing (Boyer and Marcy, 2008), canonicalizing means encoding the XML document in a single charset (UTF-8), adding default attributes, and many more changes as listed in the specification (Boyer and Marcy, 2008).
Canonicalization is useful in the context of build reproducibility as it allows to convert the output artifacts of the build process into a standard representation that is independent of the machine and environment where the build has been performed. Assume that the output artifact contains details of the operating system or the user who built the artifact. These details are not relevant to the functionality of the software and can be removed as part of canonicalization or the artifact.
2.2. Build Reproducibility
Build Reproducibility is a property of a software build process where the output artifact is bit-by-bit identical when built again, given a fixed version of source code and build dependencies, regardless of the environment (Lamb and Zacchiroli, 2022). This is shown in the equation below, where source module is built twice by builder and rebuilder to produce reference and rebuild artifacts, respectively. The build is reproducible if the reference and rebuild artifacts are identical.
Reproducible builds help prevent software supply chain attacks at the build time by ensuring that the software is not tampered with during the build process. It also allows users to verify that the executable provided to them is free of backdoors as they can build the software themselves and compare the output with the provided executable.
(1) $ReferenceArtifact\leftarrow Build(SourceModule)$
(2) $RebuildArtifact\leftarrow Rebuild(SourceModule)$
(3) $ReferenceArtifact=RebuildArtifact$
There are 2 other terms in the literature that are used interchangeably with build reproducibility - “verifiable builds” and “accountable builds” (Pohl et al., 2025). Builds are verifiable if the build process can be modified in order to make the output artifact bit-by-bit identical. Accountable builds are the same as verifiable builds, but instead of removing the differences, these differences are explained and documented so that unreproducibility can be understood upon inspection. In our paper, we stick with the term “build reproducibility” as the other two terms are extensions of build reproducibility with extra steps.
2.3. Mitigation
We claim that any cause of unreproducibility can be mitigated in three ways - fixing the build process, fixing the rebuild process, or canonicalizing the output of the rebuild process.
Fixing the build process and Fixing the rebuild process involve modifying inputs such as source code, dependencies, or build tools (e.g., build scripts and configurations) to the build or rebuild workflows in order to eliminate sources of non-determinism. The builder ensures that the build process is deterministic and easy to reproduce for rebuilder. The rebuilder tries to perform the build process as closely as possible to the original build process. However, some build tools generate non-deterministic information or there is a difference in the environment which makes these mitigation techniques not always feasible neither practical.
Canonicalizing the output of the rebuild process as defined above means removing non-deterministic information from output artifacts. Some non-deterministic information is spurious, but it is sometimes perfectly legitimate such as a cryptographic signature with one-time key (de Carné de Carnavalet and Mannan, 2014). Canonicalization is generally not done by the builder for this reason and it is more the responsibility of the rebuilder. It requires careful removal of all non-deterministic information from the output artifact. If done incorrectly, it may inadvertently remove or obscure meaningful differences in the output artifacts. This can lead to false positives, where builds that are genuinely unreproducible are incorrectly reported as reproducible. Such errors would undermine the reliability of reproducibility verification and can mask underlying issues that need to be addressed. Therefore, it is crucial to ensure that canonicalization is performed with precision, preserving all relevant differences while eliminating only the spurious, non-deterministic variations. Canonicalization can address the unreproducibility issues of past releases by removing non-deterministic information from its output artifacts. It can also help as last resort to fix unreproducibility issues when build process and rebuild process are embedding non-deterministic information which is not in control (like cryptographic signatures). Moreover, techniques related to fixing build process have also been studied in prior literature (Randrianaina et al., 2024; Ren et al., 2022; Mukherjee et al., 2021; Xiong et al., 2022; Keshani et al., 2024; Macho et al., 2021). Although we suggest these techniques as mitigation strategies, we focus deeply on the third mitigation strategy (canonicalization) due to the fact that it is not well evaluated in the literature.
3. Dataset of Unreproducible Builds
3.1. The Reproducible Central Project
The Reproducible Central project (Boutemy, 2024) is a rebuild infrastructure to verify whether Maven projects are reproducible. It specifies list of steps to build a Java project from source in a buildspec file  [^7] https://github.com/jvm-repo-rebuild/reproducible-central/blob/master/doc/BUILDSPEC.md. Then, the Reproducible Central script takes this buildspec as input to rebuild a project. It produces buildinfo and buildcompare files; buildinfo records the output of the build in terms of the artifact names, sizes, and checksums (Builds, 2024); buildcompare reports which artifacts are reproducible and which are unreproducible.
We use the Reproducible Central project dataset as a starting point to study the causes of unreproducible builds in Java. We take a snapshot of the Reproducible Central dataset on 8th October, 2024 whose commit is d280bf1. At this date, Reproducible Central contains 706 Maven projects, their corresponding build scripts, and a history of 4,956 releases. It is the largest dataset attempting rebuilds of Java projects and is also maintained by maintainers of Apache Maven.
Our goal is to curate a dataset of unreproducible released Maven modules in order to study the reasons why unreproducible builds occur. In order to curate the dataset of unreproducible released modules, we first apply 2 filters before running the rebuilds and 1 filter afterwards.
3.2. Filtering
3.2.1. Maven Projects
First, we filter out Maven projects that require manual intervention to build so that we can run the entire rebuild process automatically. This is indicated in buildspec file either by the presence of build tool or when the rebuild command is prefixed with ‘SHELL’. The rebuild script would spawn an interactive terminal when any of the above conditions are met. We exclude projects that do not have a POM file at the root directory as it also requires manual intervention to get the path to rebuild artifacts in build directory. Second, we only include projects that use Maven as their build tool (some projects use Gradle or SBT). This is because Maven is the most mature build tool for Java projects, most widely used in the Java ecosystem (Sulír and Porubän, 2016). Moreover, Maven builds success rate is higher than Gradle and Ant (Maes-Bermejo et al., 2022). Once all rebuilds are done, we exclude projects where the unreproducible artifacts are not a subset of the all artifacts released on Maven Central. This can happen when not all artifacts produced by the build process are released on Maven Central. For example, content/org.infinispan:protostream-aggregator:5.0.7.Final produces an artifact rotostream-integrationtests-5.0.7.Final-test-sources.jar which is not released on Maven Central. After applying the filters, we end up with 3,927 releases over 665 Maven projects.
3.2.2. Released Modules
Next, we need to identify the released modules to compare against. In order to map each unreproducible project to corresponding released modules, we first identify the Maven modules in each project. org.apache.aries.cdi:org.apache.aries.cdi.executable:1.1.5 is an example of a source module that does not have a corresponding URL on Maven Central. We parse the POM file located at the root directory of the project as it contains the required information about the modules. For example, jpmml-python is a Maven project hosted on GitHub  [^8] https://github.com/jpmml/jpmml-python/tree/1.2.2 which has 3 Maven modules with version 1.2.2 - org.jpmml:jpmml-python, org.jpmml:jpmml-python:1.2.2, and org.jpmml:pmml-python. Thus, jpmml-python Maven project on Reproducible Central corresponds to 3 Maven modules in our dataset. 2 out of 3 released modules in this project contain unreproducible artifacts. org.jpmml:pmml-python-testing:1.2.2 has pmml-python-testing-1.2.2-sources.jar and org.jpmml:pmml-python:1.2.2 has pmml-python-1.2.2-sources.jar. Finally, there are 4 cases where unreproducible reference and rebuild artifact are not mapped correctly. Maven module org.apache.ratis:ratis-assembly:3.1.0 and 3 more of its versions have an artifact with .pom extension. The rebuild script maps this artifact to the Jar file produced during rebuild. This is a bug in the Reproducible Central rebuild script and have reported this to the author  [^9] https://github.com/jvm-repo-rebuild/reproducible-central/issues/2164.
At the end of the pipeline, we end up with 7,961 Maven modules which have a total of 34,271 artifacts. Out of 34,271 artifacts, there are 12,283 unreproducible artifacts (35.8%).
3.3. Description of the Dataset
The dataset contains groupId, artifactId, version, and the URL of all 7,961 unreproducible released modules on Maven Central. It also contains the reference and rebuild artifact pair of all 12,283 unreproducible artifacts. It is available at https://github.com/chains-project/reproducible-central/. To our knowledge, this is the largest ever curated dataset of unreproducible artifacts in the context of Java. Based on the dataset, we can answer the following research questions:
  • RQ1 (Causes): What are the causes of unreproducible builds in Java?
  • RQ2 (Artifact Canonicalization): To what extent does artifact canonicalization using OSS-Rebuild make a Maven release reproducible?
  • RQ3 (Bytecode canonicalization): To what extent does bytecode canonicalization with jNorm mitigate unreproducible builds in Java?
4. RQ1: Taxonomy of Unreproducibility Causes
4.1. Objective
The goal of this section is to provide a comprehensive taxonomy of unreproducibility causes in Java. We first report the cause based on our analysis and then propose a mitigation strategy for each cause.
4.2. Methodology
We have a dataset of 12,283 unreproducible artifacts and for each of them we have the reference artifact from Maven Central and its rebuilt version from our own rebuild. We run diffoscope 285 on each pair of reference and rebuild artifacts to get the content diff. diffoscope is a state-of-the-art tool for computing the difference between various types of files, archives, and directories  [^10] https://diffoscope.org/. This gives us a total of 12,283 diffoscope files. The median size of diffoscope files is 102 KB, the maximum is 3.40 GB, and the minimum is 0.38 KB.
Given the scale of the dataset, we first analyze a random sample of the diffoscope files to understand the reasons of unreproducibility. There are two attributes in the diffoscope file that are of interest to us - source1 and unified_diff, both of them indicating a reproducibility problem. The source1 attribute either contains the name of the file or the command that is run to get the textual output of the file. For example, if source1 includes cyclonedx.json, we can infer that the difference is in the Software Bill of Materials (SBOM). Another example is if source1 is javap -verbose -constants -s -l -private {}, it means that the Java class file is being compared and the reason of unreproducibility is due to differences in the bytecode. The unified\_diff attribute contains the diff of the file in the unified format, for textual files. We then write regular expressions to match these attributes to categories. For example, if regex
[…]
is matched, we infer that the reason of unreproducibility is due to timestamps in the file. We define regular expressions, which we match against all 12,283 diffoscope files.
Finally, we manually go through samples that remain unmatched to the regular expressions and create expressions to map them too. This refines the different root causes of unreproducibility. Then, we report on mitigation strategies from the literature that could fix the reproducibility problem.
4.3. Results
Reason for Unreproducibility Root Cause of Unreproducibility Novelty Example Main Mitigation
Build Manifests Environment - Built-By attribute Canonicalization by rebuilder
Rebuild Process - Build-Jdk attribute Fix rebuild process
Non-deterministic configuration Embedded branch names Fix build process
SBOM Java Vendor Different checksum algorithms Fix rebuild process
Custom release configuration Releasing a subset of artifacts Fix build process
Non-deterministic information Serial Number Canonicalization by rebuilder
Filesystem Environment - Permissions Canonicalization by rebuilder
Custom release configuration Generated binaries are not included Fix rebuild process
JVM Bytecode JDK Version - Ordering of constant pool entries Fix rebuild process
Embedded data Git branch embedded Fix build process
Build time generated code - Lambda functions Canonicalization by rebuilder
Versioning Properties - - Number of tags embedded in Jar Canonicalization by rebuilder
Timestamps - - Embedded in shell script in Jar files Fix build process
Table 1. The taxonomy of unreproducibility causes based on the analysis of 12,283 unreproducible artifacts. We are the first to state the most appropriate mitigation strategy.
In this section, we present a comprehensive taxonomy of unreproducibility in Java, based on our large-scale, systematic protocol presented in Figure 1. Table 1 summarizes the taxonomy, before deep-diving into each of them. The first column lists the reason for unreproducibility. The second column further breaks down the reason into root causes. The third column indicates if the root cause is novel and has not been reported in the literature before, to our knowledge. The fourth column gives an example of the root cause that appears in our dataset. Finally, the last column lists the main mitigation strategy to fix the unreproducibility, one of the three presented in subsection 2.3.
For example, consider the row about JVM bytecode, where unreproducibility is due to changes in the JVM bytecode between the reference and rebuild artifacts. One of its root cause is due to embedded data in the bytecode, this cause has not been discussed before. The main mitigation strategy for this problem is to fix the build process by preventing build tools to embed branch names in the JVM bytecode. Branch names are mutable and can change over time. Hence embedding them in the bytecode can lead to difference in subsequent builds. The other two mitigation strategy would also work, but they have cons - 1) third-party rebuilder would have to debug the build process to understand what build tool is responsible for embedding and 2) branch names have no conventional names and hence canonicalization would risk removing meaningful differences in the bytecode.
Next, We describe each reason and its root cause in details in the following subsections.
4.3.1. Unreproducible Build Manifests
We consider two manifest files in our analysis - MANIFEST.MF and pom.properties. The MANIFEST.MF file is a mandatory metadata file in a JAR file. It potentially contains information on the entry point of the JAR file, signatures of the classes, and details of the version control system  [^11] https://docs.oracle.com/javase/tutorial/deployment/jar/manifestindex.html. pom.properties is an automatically generated file by Maven that contains the group ID, artifact ID, and version of the Maven module.
The first set of differences in manifests is due to a change in environment. For example, the Built-By attribute in the MANIFEST.MF file is sometimes updated to the username of the user who built the JAR file. Mitigation: This non-deterministic information should be stripped out from the JAR file by rebuilder before comparison. See section A.1, section A.4, section A.5, and section B.3 for more attributes that follow the same problem and solution.
The next set of differences in manifests is due to differences in the script that configures the rebuild. For example, the Build-Jdk attribute depends upon the version of JDK used to build the JAR file. Mitigation: All environment specifications must be exactly specified in the the build script by rebuilder. Ideally, the build script file extracts information from the MANIFEST to be reproduced. See section A.2, section A.3, section A.6, section A.7, and section A.8 for concrete examples.
Finally, there are reasons of differences in the manifests MANIFEST.MF and pom.properties that are due to non-deterministic behavior in the build script of the builder. For example, build manifests pom.properties and MANIFEST.MF can differ in order of the attributes if they are not sorted. Mitigation: The builder should depend upon on plugins that generate sorted data  [^12] https://github.com/apache/maven-archiver/commit/763a940540eefad74f9ba73cb5eed288dc4e639d, should not embed non-deterministic data (eg branch names) in the artifact, should ignore timestamp, and stay up to date with the latest version of plugins. section A.9, section A.10, section A.11, section B.1, and section B.2 describe these reasons and solutions in detail.
4.3.2. Unreproducible Software Bill of Materials
Software bill of material (SBOM) is a formal, machine-readable inventory of software components, metadata about those components, and the hierarchical relationships between them  [^13] https://www.ntia.gov/sites/default/files/publications/sbom_at_a_glance_apr2021_0.pdf. Its goal is to enable transparency of the software supply chain of an application, for its consumers. In the case of Java, the best practice is to push SBOMs as part of the Maven release (Gamage et al., 2025). Our experiment finds that this poses reproduciblity challenges.
The first set of changes are due to incorrect Java vendor used by the rebuilder. For example, old JDK versions do not have newer hash algorithms like SHA-3. However, vendors of some JDKs backport these algorithms and if builder uses a JDK with backported algorithms, the resulting SBOM will have hashes computed using more algorithms. If rebuilder does not use the same JDK, there will be differences in number of hash algorithms used. Mitigation: The rebuilder should use the same JDK vendor and version as the builder. See section C.1 for more details.
The next set of changes are due to a tailored release configuration by the builder. For example, the builder can choose to release only a subset of the artifacts produced by the build process which can lead to less components in the released SBOM. Mitigation: The builder should use a conventional build process - which is to streamline the build process only using Maven commands. This helps the rebuilder to reproduce the build process exactly without knowing all ad hoc scripting in the build. See section C.2 and section C.3 for concrete examples.
Finally, the last set of changes are due to non-deterministic information in the SBOM. CycloneDX attributes like timestamp and serialNumber are non-deterministic and hence they differ in each build. Mitigation: There are two possible mitigations. First, the builder can generate the values of these attributes so that they are deterministic. Second, the rebuilder can strip out these attributes from the SBOM before comparing the outputs. See section C.5, section C.4, and section C.6 which outlines the implementation on how to either generate deterministic values or strip out the non-deterministic values.
Reproducible SBOMs is an open problem in the field, we are the first to report about it.
4.3.3. Unreproducible Filesystem
We observe unreproducible Maven releases due to presence or absence of files, changes in the absolute file paths and their related metadata such as permissions, ownership, size and type. We classify them into 2 reasons below.
The first set of changes are due to an influence from the environment. For example, the permissions of the file can vary depending on the umask of the environment of rebuilder section D.2. The ownership and timestamp of the file can also vary depending who and when the artifact is built section D.3. File sizes can vary depending on the file system used to build the artifact section D.6. The names of the files can vary if absolute paths are used to refer to the files in artifact section D.7. These names can be of generated files or they could be embedded in JVM bytecode and configuration files which makes the artifact unreproducible section D.8. Finally, the order of files in Jar file section G.5 can vary as the order depends upon the filesystem layout  [^14] https://reproducible-builds.org/docs/archives/. Mitigation: The build script should either set a fixed value or strip this information from reference and rebuild artifacts before comparison.
Finally, there are some causes due to differences in the build process by the builder and rebuilder. For example, some binaries are not included in the rebuild version as they are not generated by the rebuilder build process. Mitigation: The rebuilder process should replicate the build process of the builder exactly. For the above example, the rebuilder should incorporate commands to generate the missing binaries. See section D.1 and section D.5 for more concrete examples.
4.3.4. Unreproducible JVM bytecode
We observe 898 unreproducible Maven releases due to changes in the JVM bytecode when rebuilding. The reasons are as follows.
The first set of changes are due to the use of different versions of the JDK or JDK flags used in the build and rebuild process. This differences are due to core compilation changes such as refactoring, ordering changes in the bytecode (specifically methods, fields, static initializers, entries in the constant pool, and array type values in annotation section G.1), optimization flags, debug flags. For example, java.nio.ByteBuffer.flip() returns a different type in JDK 8 and JDK 11. For io.dropwizard.metrics:metrics-collectd:4.1.20 rebuilt with JDK 11, all flip calls return java.nio.ByteBuffer which is different from the reference version built with JDK 8 where the return type is java.nio.Buffer.  [^15] This is reported in the bug tracking system of OpenJDK (Gafter, 2002) which shows that this change is introduced in JDK 9. See section E.1, section E.2, and section E.3 for more examples. Mitigation: The rebuilder should use the same JDK version and flags as the builder. We propose that the builder should always embed the precise vendor and version of JDK used to build the artifact in the MANIFEST file under the attribute Build-Jdk. This way, the rebuilder can use the same version and vendor of JDK to rebuild the artifact by reading the information instead of guessing it. If the builder is using a custom JDK as done in Google Guava (Moore, 2023), this should be made available to the rebuilder for verification.
The next set of changes are due to embedding information related to environment. We observe that the bytecode has git branch names, user name of the builder, and absolute paths. For example, org.apache.hive:hive-exec:4.0.0-alpha-2 embeds git properties and timestamps in package-info.class file using a shell script that adds a Java annotation to package-info.java. The Java compiler also embeds the exact Java version used to compile the source code in the bytecode of module-info. However, this hinders reproducibility even if the patch version of the JDK is different. Mitigation: The builder should avoid embedding environment specific information in the bytecode. This information changes with each rebuilder and hence the bytecode will be different making the artifact unreproducible. See section E.4 for example of a release where it happens.
Finally, the last set of changes are due to build time code generation or modification. The Java compiler produces synthetic code during build in order to handle access to type members (Rimenti, 2018). The compiler also names lambda functions (Evans, 2020) and anonymous classes (Maverik, 2016) which are not present in the source code. In addition to these capabilities, Maven provides plugins to generate or modify code during the build process. All of this generated code can be different across different builds and hence the build is unreproducible. For example, we observe a difference in names of lambda function in org.apache.helix:zookeeper-api:1.0.3. Another example is org.apache.hive:kafka-handler:4.0.0-alpha-2 which uses maven-shade-plugin to rename package names from kafkaesqueesque to kafkaesque but only when a Maven profile is activated. Mitigation: The names of the synthetic code can be canonicalized by the rebuilder before comparison. Bytecode canonicalization (Sharma et al., 2024) is a technique to strip out non-deterministic information from the bytecode. Differences caused by Maven plugins can be mitigated by running the same plugins as in the original build process. Maven plugins can be configured to run only for certain build profiles. The rebuilder should document the build profiles used by the builder to reproduce the build process. See section E.5 for more examples.
4.3.5. Unreproducible Versioning Properties
Git properties are embedded in Jar files in git.json or git.properties files. Maven plugin git-commit-id-maven-plugin  [^16] https://github.com/git-commit-id/git-commit-id-maven-plugin is commonly used to do that. These properties include timezone of commit, remote URL, number of git tags in repository, name of build host, time of build, name of builder, email of builder, branch name, number of commits, local branch name, and total commits in the repository. All of them sometimes diverge. For example, value of git.tags vary for Maven release org.apache.drill:drill-opentsdb-storage:1.21.0 due to additional number of tags in the repository during rebuild. Mitigation: The rebuilder should strip out the git properties from the Jar file before comparison or avoid comparison of the Git properties. Another solution is to fix the values for each git property. For example, total commits should be calculated up to the commit that is used to build the artifact. See issue on our repository  [^17] https://github.com/chains-project/reproducible-central/issues/19 to see more examples of differences in the git properties.
4.3.6. Unreproducible Timestamps
Timestamps are the most common cause of unreproducibility in our dataset as we observe them embedded in multiple ways spread across artifacts in Maven release. They are a widely known cause of unreproducibility (Bajaj et al., 2023) (Goswami et al., 2020) (Benedetti et al., 2025).
In Maven, one solution exists to prevent timestamps from breaking reproducibility. The Maven POM file can be configured to set the timestamp to a fixed value using property project.build.outputTimestamp  [^18] https://maven.apache.org/guides/mini/guide-reproducible-builds.html. The value of this property is then used by Maven plugins like maven-jar-plugin to set the timestamp in the JAR file.
However, we observe that there are 10 other ways timestamps can be embedded in the JAR file. In our dataset, we observe that: 1) they can be embedded in the properties file section F.1, 2) generated documentation section F.2, 3) shell scripts section F.3, 4) executable binaries section F.4, 5) software bill of materials section F.5, 6) JVM bytecode section F.6, 7) file metadata section F.7, 8) MANIFEST.MF section F.8, and 9) NOTICE section F.9. 10) servlets created by Jasper JSP compiler section F.10. Mitigation: The solution proposed by Maven is partial, more engineering is needed to support project.build.outputTimestamp in all Maven plugins that create the above mentioned files.
4.3.7. Miscellaneous Unreproducibility Reasons
We report reasons of reproducibility that are too specific to be categorized in the above sections.
The first reason is due to the use of different build tools. For example, the Maven build tool is available as standalone tool and it also integrated into the Eclipse IDE. They embed information differently. m2e.projectLocation and m2e.projectName is embedded in pom.properties of org.spdx:spdx-maven-plugin:0.7.0 and org.owasp.antisamy:antisamy:1.7.3 section B.3. Even if the build tool is same, there are differences in rebuild due to usage of versions ranges of a plugin. For example, org.apache.aries:org.apache.aries.jax.rs.whiteboard:2.0.2 uses [1.2.5,) as the version range for flatten-maven-plugin in the POM file. This causes difference in the ouput of the plugin. Mitigation: The rebuilder should use the same build tool and its precise version as the builder. Moreover, the builder should avoid using version ranges for plugins in the POM file.
The next reason is due to a difference in the operating system. The reference version of org.apache.helix:helix-front:1.0.3 produces a Mach-O object file format (Mac) which is different from the ELF binary format produced by the rebuilt version (Linux). See https://github.com/chains-project/reproducible-central/issues/20#issuecomment-2583944777 for detailed diff. Mitigation: The builder should ensure that 1) the OS is specified in the build specification or 2) that the binary format is independent of the operating system used to build the artifact.
Usage of Universally Unique Identifier (UUID) is another reason of unreproducibility. We observe that Maven release org.apache.karaf.examples:karaf-docker-example-static-dist:4.4.0 uses UUID to generate the configuration files. Mitigation: The builder should use seed values to generate the UUID so that the UUID can be reproduced.
Maven releases org.apache.helix:helix-front:1.0.3, io.wcm:io.wcm.caconfig.editor:1.9.0, and org.apache.isis.viewer:isis-viewer-restfulobjects-viewer:2.0.0-M7 embed bundled JavaScript files in the Jar file. These files are subject to changes in variable names and new lines. Mitigation: The changes in variable names can be fixed by pinning the version of the bundler. For example, Goswami et al. (Goswami et al., 2020) show that UglifyJS can produce output with different variable names for the same input if the version of UglifyJS is different. The difference in new lines can be fixed by the rebuilder by canonicalizing them.
There is one instance where microbenchmarking results are embedded in the Jar file section G.4. Mitigation: The builder should not embed microbenchmarking results in the Jar file as they are solely for the purpose of improving performance (Costa et al., 2021) of the application and not needed for the end user.
Finally, there are some indices files created by tools like Jandex  [^19] https://smallrye.io/jandex/jandex/3.2.1/index.html - a space efficient Java class file indexer, EFXToolkit  [^20] https://github.com/OP-TED/efx-toolkit-java - a toolkit for expressing eForms business rules, and SpotlessFormatter  [^21] https://github.com/diffplug/spotless - a code formatter for Java. Mitigation: The builder should not embed these index files in the artifact as they are not required for the end user and they serve only as cache files for plugins that use them.
4.3.8. Mysteries
We observe changes in NOTICE, POM, and other configuration files that are not due to any of the above reasons and thus, we are unable to propose a mitigation strategy for them. Since these reasons of unreproducibility are so diverse, we discuss them one by one in appendix H.
{mdframed}
Answer to RQ1: “What are the causes of unreproducible builds in Java?”
We identify 6 main reasons of unreproducibility in Java and their mitigation: 1) build manifests, 2) software bill of materials, 3) filesystem, 4) JVM bytecode, 5) versioning properties, and 6) timestamps. They represent the key challenges that researchers will address. For practitioners, they provide clear guidelines to diagnose and mitigate their unreproducible builds in Java.
5. RQ2: Unreproducibility Mitigation with Artifact Canonicalization
5.1. Objective
The goal of this research question is to investigate if artifact canonicalization can make the build reproducible.
5.2. Methodology
Artifact canonicalization means that the entire artifact is transformed by removing non-deterministic and spurious changes, especially in metadata. This helps downstream build tools to interpret the artifact in a consistent way. Multiple reasons can prevent the reproducibility of an artifact( section 4), such as the presence of non-deterministic information in files like MANIFEST.MF, pom.properties, and git.properties files. Moreover, the artifact may contain non-deterministic information in the form of timestamps or file order. Artifact canonicalization is a process of transforming the artifact into a representation that is independent of these spurious changes. For example, the pom.properties file is a file generated by Maven that records the group ID, artifact ID, version, and timestamp of the build in the JAR file. Artifact canonicalization can also remove the timestamp and set a fixed order to the declaration of properties.
OSS-Rebuild  [^22] https://github.com/google/oss-rebuild/commit/4ef4c013fe6903cda40a9ee4244e3b65b5834325 is a tool by Google to canonicalize software artifacts in order to improve reproducibility. In OSS-Rebuild, “canonicalizing” is called “stabilizing”. It is capable of canonicalizing tarballs, ZIP files, and GZIP files. For example, it canonicalizes file order, modification timestamp, compression algorithm, encoding, and other miscellaneous attributes from a ZIP archive.
Chains-Rebuild  [^23] https://github.com/chains-project/chains-rebuild/commit/6dd67d5c7ac4db112f3419b5132d8f80a22cbe65 is our own fork of OSS-Rebuild, with crucial improvements dedicated to Maven reproducibility, which is unfortunately missing in the upstream project. We add support for canonicalizing build manifests and embedded versioning properties in JAR files as suggested in taxonomy (Table 1). For example, the Built-By attribute in the MANIFEST.MF file is sometimes updated to the username of the user who built the JAR file, a spurious change which is removed by Chains-Rebuild. For versioning properties, we remove the git.properties files embedded in the JAR file. Finally, we remove the pom.properties file.
We run OSS-Rebuild and Chains-Rebuild on the 12,283 reference and rebuild artifact pairs in order to analyze how effectively they make the build reproducible, that is when artifacts are bit-by-bit identical after canonicalization as shown below.
$canonicalize(ReferenceArtifact)=canonicalize(RebuildArtifact)$
5.3. Results
Tool Successful Failed
Canonicalization Canonicalization
OSS-Rebuild (4ef4c01) 1165 (9.48%) 11118 (90.52%)
Chains-Rebuild (6dd67d5) 3036 (24.72%) 9247 (75.28%)
Table 2. Effectiveness of OSS-Rebuild and Chains-Rebuild on mitigating 12,283 unreproducible artifacts via canonicalization.
Table 2 shows the results of OSS-Rebuild and Chains-Rebuild on 12,283 unreproducible artifacts. The first column is the name of the tool and its commit hash we use to canonicalize the artifacts. The second column shows the number of artifact pairs that are successfully canonicalized by OSS-Rebuild and Chains-Rebuild. The third column shows the number of artifact pairs that are not properly canonicalized.
OSS-Rebuild, with generic archive stabilizers, is little effective, with only 9.48% of the artifacts becoming reproducible. Chains-Rebuild is significantly more effective, with 2.5x more success, thanks to dedicated features for canonicalizing JAR files. Recall that, we add support for canonicalizing MANIFEST, embedded versioning properties, and pom.properties files in JAR files. For example, artifact slf4j-ext-2.0.6.jar of Maven release org.slf4j:slf4j-ext:2.0.6 is successfully canonicalized by Chains-Rebuild. Originally, there are differences in zip metadata and MANIFEST.MF. OSS-Rebuild only canonicalizes the zip metadata. Chains-Rebuild takes a step further and canonicalizes the MANIFEST.MF file as well. It fixes the ordering of values in MANIFEST.MF under the Export-Package attribute. As a minified example, the fix changes Export-Package:org.slf4j.ext;version="2.0.6",org.slf4j.agent;version="2.0.6" to Export-Package:org.slf4j.agent. Here arranging the order of the values in Export-Package attribute leads to a canonicalized output and a reproducible build.
Overall, this experiment is evidence that artifact canonicalization is a valid solution to mitigate some unreproducible builds. However, it is only a partial workaround and does not replace the mitigation fixing build and rebuild process (see subsection 2.3). As shown in Table 1, there are many other causes of unreproducibility that can be fixed by canonicalization. For example, non-deterministic information in SBOMs and environment influences in files can be fixed by canonicalization. This includes SBOM, generated JVM bytecode, and ordering of sections in JVM bytecode. We leave this as future work as it is purely an engineering task.
{mdframed}
Answer to RQ2: “To what extent does artifact canonicalization using OSS-Rebuild make a Maven release reproducible?”
Our publicly available prototype Chains-Rebuild can canonicalize (3036) 24.71% of the artifacts. Our study is the first large-scale experiment on artifact canonicalization, demonstrating that it is a promising mitigation of build unreproducibility in Maven.
6. RQ3: Unreproducibility Mitigation with Bytecode Canonicalization
6.1. Objective
In section 4, we observe that changes in JVM bytecode is one of the causes of unreproducible builds. In this research question, we investigate to what extent these changes can be mitigated by using bytecode canonicalization.
6.2. Methodology
Bytecode canonicalization is a process of transforming the bytecode of a program into a representation that is independent of specific implementation details inserted by the compiler. This is either done by converting the bytecode to an intermediate form or by rewriting the bytecode. For example, subtraction of two integers is implemented as addition of a positive and negative integer in Java 6 and higher versions. Meanwhile, in Java 5 and lower versions, it is implemented as subtraction of two integers. Hence, if the program is compiled with Java 5 and 6, the two bytecode representations differ even though the program is semantically the same - performing subtraction of two integers. Canonicalization removes these differences by converting the bytecode to an intermediate form which is free from the implementation details that vary across different versions of the compiler.
We consider the state-of-the-art Java bytecode canonicalization tool by Schott et al. (Schott et al., 2024), called jNorm. jNorm’s primary goal is to remove the parts from Java bytecode that may differ due to changes in JDK version and output the canonicalized version of the bytecode. jNorm only removes the differences that do not change the semantic behavior of the artifact. Note that they refer to “canonicalization” as “normalization” in their tool.
We run jNorm version 1.0.0 on the 898 reference rebuild artifact pairs that have differences in the bytecode. For maximum canonicalization, we run the tool with all the command line options as listed in the README 242424jNorm tool - https://github.com/stschott/jnorm-tool/tree/cec4645c5c9b52f73c347349bf14945b0eb55c87 of the repository and tool’s help output. As a result of canonicalizing the artifacts, jNorm produces files in the Jimple (Raja Vallée-Rai and Laurie J. Hendren, 1998) format. We save the canonicalized Jimple files for both the reference and the rebuild artifacts. There are cases where jNorm crashes and canonicalization fails, which we consider as a failure. We do not analyze them further as they are bugs in tool, but we upload the logs for future debugging. We run the Linux diff tool over the canonicalized pairs of Jimple files. If there is no difference, it means that canonicalization is a success, for reproducibility. In other words, successful canonicalization by jNorm produces two identical artifacts. We report the number of cases where the diff between the pair is empty. Finally, we analyze the cases where jNorm is unable to fully canonicalize the bytecode in order to understand its fundamental limitations and identify required future work.
6.3. Results
Out of 898 artifacts, jNorm’s execution leads to the following distribution: it successfully canonicalizes 267 (29.7%) artifacts which means it produces two identical artifacts; it leaves 478 (53.2%) artifacts with a diff after canonicalization; and there are 153 (17.1%) crashes, which are ignored.
Our results are in contrast with the results of Schott et al. (Schott et al., 2024) where the authors report that jNorm is able to canonicalize up to 99% of pairs of Java classfiles when JDK version differs. However, our success ratio is lower (29.7%). This can be attributed to the different methodology used to generate the dataset. jNorm is evaluated on Java classfiles built on the same machine but varying JDK versions. In our case, the reference and rebuild artifacts are built on different machines and based on our results in section 4, we know that the build process is different.
Next, we study the reasons why jNorm fails to produce identical artifacts when run on 478 cases. We categorize them based on the features in the bytecode that are not properly canonicalized by jNorm. Note that some unreproducible artifacts may fall into multiple categories as it is possible jNorm is unable to canonicalize multiple problems in the same artifact.
Structural Change Limitation
Recall from subsubsection 4.3.4 that the order of methods, fields, static initializers, entries in the constant pool, and array type values in annotation sometimes differ between the reference and rebuilt artifacts. We find that jNorm is unable to canonicalize most of those problems: the order of fields, methods, static initializers, array type values in annotation, removal of implicit modifier from fields. The only mitigated problem by jNorm is the constant pool order problem, because the Jimple format abstracts away the constant pool (Vallée-Rai et al., 1999). We now discuss concrete examples of failed canonicalization:
  • org.apache.shiro:shiro-aspectj:1.11.0 is not canonicalized due to the changed order of static initializers in the bytecode.
  • antisamy-1.7.3.jar has a difference in implicit visibility modifier between the reference and rebuild artifact, not handled by jNorm.
  • org.apache.helix:zookeeper-api:1.0.3 has one unreproducible artifact which is due to the naming of lambda functions in the bytecode. Since the names of lambda functions are suffixed by integers determined non-deterministically by the compiler, jNorm is unable to canonicalize them.
Control Flow Limitation
There are two cases of this category where jNorm is unable to canonicalize changes in control flow.
  • The first case is change in control flow of if statements as illustrated in LABEL:lst:javap-diff and LABEL:lst:jnorm-diff. eu.maveniverse.maven.mima:2.4.2:standalone-static-uber have changes in the control flow of if conditions. The reference checks for the negative condition of the if statement while the rebuild checks for the positive condition first, because the two builds use different compilers. This is shown in the diff generated by diffoscope in LABEL:lst:javap-diff. jNorm does not canonicalize control flow there is still a diff in Jimple format, as seen in LABEL:lst:jnorm-diff.
    1  - 99: ifne 106
    2 - 102: iconst_1
    3 - 103: goto 107
    4 - 106: iconst_0
    5 - 107: ireturn
    6 + 99: ifeq 104
    7 + 102: iconst_0
    8 + 103: ireturn
    9 + 104: iconst_1
    10 + 105: ireturn
    Listing 1: javap diff for control flow difference.
    1  -if v != 0 goto label;
    2 -v = 1;
    3 -goto label;
    4 -label:
    5 -v = 0;
    6 +if v == 0 goto label;
    7 +return 0;
    8label:
    9 -return v;
    10 +return 1;
    Listing 2: jNorm diff for control flow differnce.
  • Second, Maven release io.fabric8:kubernetes-model-apiextensions:6.4.0 contains generated code where the order of if-else blocks is different across different builds. This is not an issue with jNorm, this is an issue with the unreproducibility of the code generator employed during build.
Embedded Data Limitation
In section 4, we discussed that absolute file paths, timestamps, Java version, and project version are embedded in the bytecode. jNorm is able to canonicalize the Java version and project version as it deletes the module-info.class file that stores this information. However, it is unable to canonicalize absolute file paths and timestamps252525jNorm is able to canonicalize timestamp if they are declared in a static final field..
Optimization Limitation
We find 5 cases of optimizations or de-optimizations that are not canonicalized by jNorm. We present the different features in Java that are not canonicalized by jNorm. For a more detailed discussion, refer to the  Appendix I.
  • JDK 17 creates a lookup table to store the values of the enum while JDK 22 simply stores it based on the order of the enum values.
  • JDK 17 uses the invokevirtual bytecode instruction when handling invocations to toString, equals, and getClass, while JDK 21 uses invokeinterface. This is due to a change proposed in JDK 18  [^26] https://github.com/openjdk/jdk/pull/5165.
  • String concatenation expressions where one of the operands requires a cast to String.
  • Implementation of try and try-with-resource statements and finally clauses vary between Java 5, 8, and 11.
  • Reference to outer class from inner class is handled differently in patch versions of Java 17.
We do not delegate fixing all of these problems to jNorm. Recall that we mentioned in section 4 that rebuilder can either fix the build process or canonicalize the output in order to mitigate the unreproducibility. jNorm can fix problems in bytecode structure, control flow, and embedded data. However, unreproducible optimization and de-optimization can be mitigated by fixing the Java version in the build process.
We note that multiple canonicalization tools can be used in conjunction. For instance, Chains-Rebuild and jNorm can be used together to canonicalize the JVM bytecode and the JAR files of a Maven artifact. Since Chains-Rebuild can canonicalize 3,036 artifacts and jNorm can canonicalize 267 artifacts, we can combine the 26.89% (3,303 / 12,283). Out of 3,303 artifacts, 2,578 artifacts are canonicalized such that 2,263 / 7,961 (28.43%) releases become fully reproducible.
{mdframed}
Answer to RQ3: “To what extent does bytecode canonicalization with jNorm mitigate unreproducible builds in Java?”
jNorm is able to canonicalize 267 (29.7%) out of 898 bytecode artifacts. However, there are 478 (53.2%) artifacts that are still unreproducible after canonicalization which calls for improvements. We present 4 limitation of bytecode canonicalization related to structural change, control flow, embedded data, and optimization. Finally, our experiments demonstrate that combining artifact and bytecode canonicalization successfully mitigates unreproducibility in 26.89% (3,303 / 12,283) artifacts.
7. Related Work
7.1. Reproducibility in Applications
Reproducibility is a key feature in state of the art systems. One of the first software to become reproducible was Bitcoin  [^27] https://github.com/bitcoin/bitcoin ensuring that none of the nodes in the Bitcoin peer-to-peer network contains backdoors. Tor Browser [^28] https://gitlab.torproject.org/tpo/core/tor, a web browser that anonymizes the user’s web traffic, is also reproducible since 2013 (Levsen et al., 2025). Tails, an operating system that protects against surveillance and censorship, also announced in 2017 that their ISO images are reproducible  [^29] https://tails.net/news/reproducible_Tails/. Both the Tor Browser and Tails operating system emphasize that Reproducible Builds are necessary for software geared towards privacy and security.
The Reproducible Builds  [^30] https://reproducible-builds.org/ project works on the reproducibility of coreboot  [^31] https://tests.reproducible-builds.org/coreboot/coreboot.html - an open-source firmware for x86 and ARM systems, Debian  [^32] https://tests.reproducible-builds.org/debian/reproducible.html - a Linux distribution, FreeBSD  [^33] https://tests.reproducible-builds.org/freebsd/freebsd.html and NetBSD - a UNIX based operating systems, and OpenWrt  [^34] https://tests.reproducible-builds.org/openwrt/openwrt.html - operating system for embedded devices.
In the Java ecosystem, Eclipse Temurin, a vendor for Java Developement Kit, provides reproducible builds for their JDK (Leonard, 2024). The build of Apache Maven is also reproducible  [^35] https://github.com/jvm-repo-rebuild/reproducible-central/blob/master/content/org/apache/maven/maven/README.md. Our work aims to enable ever more Java software applications to become reproducible by providing comprehensive and effective mitigation strategies.
7.2. Verifying Reproducibility
In the literature, we have found three ways to check reproducibility. First approach is to build once and compare the generated artifacts with the artifacts on package registry. Second is to build twice with variations in environment and compare the generated artifacts. Finally, third approach is to verify that the projects stand the test of time by ensuring that they build successfully over time.
Comparison between reference and rebuild artifacts
One way to verify build reproducibility is to build once and compare the generated artifacts with the artifacts on package registry. This is what we do in this paper.
Outside the Java ecosysem, Malka et al. (Malka et al., 2025) investigate reproducibility in functional package manager Nix. The paper proves that functional package managers help reproduciblity by showing 91% of the packages in Nix are reproducible. Similarly, Bajaj et al. (Bajaj et al., 2023) investigate reproduciblity in Debian and Arch Linux packages. Apart from uncovering the causes of unreproducibility, they also do a survival analysis of packages to see how long it takes to fix the unreproducibility and how long the package stays reproducible. They show that Arch Linux pacakges become reproducible sooner than Debian packages. However, a Debian package is likely to stay reproducible for longer before it becomes unreproducible again. Moritz (Moritz, 2023) proposes a decentralized protocol on Hyperledger Fabric to verify reproducibility. The protocol involves comparing the locally generated buildinfo file with the one on Debian registry. Linderud (Linderud, 2019) also proposes a protocol to verify buildinfo files, but a layer of transparency log is added in order to ensure reproducibility status for specific builds are immutable. Drexel et al. (Drexel et al., 2024) contributes with an independent rebuilder that verifies reproducibility of Arch Linux packages based on this approach. Vu et al. (Vu et al., 2021) focus on finding the most common files and APIs that differ between source repository and Python wheel published on PyPI. Goswami et al. (Goswami et al., 2020) investigate reproducibility in NPM packages. They compare the minified artifacts generated by JavaScript bundlers to report the reasons of unreproducibility. Although there are overlap in taxonomies reported by these works, our analysis focuses on Java ecosystem and proposes novel reasons for unreproducibility. For example, variations in MANIFEST, JVM bytecode, and versioning properties are first seen in our work. Even for reasons such as timestamps, we are the first to report 9 ways that it can be embedded in a Maven release.
Some works analyze single projects rather than the entire ecosystems. Shi et al. (Shi et al., 2022) investigate reproducibility for 4 commercial systems running in Huawei and CentOS - a production grade Linux operating system. Pöll et al. (Pöll and Roland, 2022) discover reproducibility issues while rebuilding Android 5 to 12 images. Carné de Carnavalet et al. (de Carné de Carnavalet and Mannan, 2014) investigate reproducibility for 16 versions of the TrueCrypt encryption tool. Finally, Andersson (Andersson, 2024) investigates reproducibility for Go Ethereum binary. None of the subjects in these related work are Java based. The work of Pöll et al. (Pöll and Roland, 2022) is Java based, but Android images do not compile to JVM bytecode like traditional Java applications. Hence, ours paper is the first one to investigate reproducibility of normal Java applications at scale.
The most closely related work here is the work by Xiong et al. (Xiong et al., 2022) that compares 59 Maven projects on Maven Central and internal registry of Huawei with the artifacts generated by the build process. They discover build reproducibility problems due to environment, JDK, multithreading, and other tools. Our work differs in the scale of the dataset (ours is 4$\times$ larger) and the taxonomy of the root causes of unreproducibility. Our taxonomy has more categories than the one proposed by Xiong et al. incl. new essential categories such as variations in filesystem and software bill of materials.
Comparison between two subsequent builds
Benedetti et al. (Benedetti et al., 2025) verify the reproducibility of 12,000 Ruby, PyPI, and Maven packages using this approach. They keep all inputs same except the environment which is changed for each build using reprotest  [^36] https://salsa.debian.org/reproducible-builds/reprotest. reprotest can change locale, timezone, and number of CPUs for example. Ren et al. (Ren et al., 2018, 2022) and Leija et al. (Navarro Leija et al., 2020) propose a technique to localize and fix the root cause of unreproducibility in Java artifacts. They build the source files and dependencies for Debian packages twice and compare the artifacts using diffoscope  [^37] https://diffoscope.org/. Lagnöhed (Lagnöhed, 2024) propose a tool rmake built on top of make with support for building projects twice and comparing the artifacts. These works differ from our work in the approach they take to verify reproducibility. We claim that the first approach uncovers more reasons of unreproducibility as it compares artifacts generated by build process which have maximum differences in environment. Carné de Carnavalet et al. (de Carné de Carnavalet and Mannan, 2014) also emphasizes that there should be maximum difference between the environments to detect unreproducibility.
Build Outcome Analysis
Build success is a prerequisite to study reproducibility. Maes-Bermejo et al. (Maes-Bermejo et al., 2022), Hassan et al. (Hassan et al., 2017), (Sulír and Porubän, 2016) and Yasi et al. (Yasi and Qin, 2022) show that only 40-60% of the Maven projects in their respective datasets build successfully. This indicates that around half of the projects don’t even generate artifacts as their build process fails. All of these works are different from ours as they focus on the success of the build process only and do not look at the generated artifacts. Our work focuses on artifacts to identify the reasons of unreproducibility. Technically, Maes-Bermejo et al. (Maes-Bermejo et al., 2022) use mvncleancompile-X to build all the Maven projects. In our work, our build commands in the dataset are tailored to each project so that the build process succeeds.
7.3. Fixing Unreproducibility
We now review the related work on fixing causes of unreproducibility.
Ren et al. (Ren et al., 2022, 2019, 2018) propose techniques to localize and fix unreproducibility in Debian packages. They localize the file in the Debian package that causes unreproducibility and then suggest a patch based on a history of patches for fixing unreproducible builds. This approach is specific for Debian packages and does not apply to Java artifacts. Moreover, their approach to detect unreproducibility is different from ours as mentioned in section 7.2.
Mukherjee et al. (Mukherjee et al., 2021) propose a technique to fix unreproducible Python builds caused by dependency errors. They report that 71% of Python builds in their dataset fail; they fix these builds by analyzing the logs. This work is only comparess success of builds, not artifacts, and it is focussed on the Python ecosystem.
Xiong et al. (Xiong et al., 2022) classifies fixing unreproducibility into 3 categories - 1) remediation by patching source code, third-party dependency, or build script 2) controlling which means interception of non-deterministic build instructions and returning pre-defined values for them 3) interpretation which means documenting the reason for unreproducibility. The authors perform these fixes automatically using a tool called JavaBEPFix which is not available to the public. Hence, we cannot evaluate the effectiveness of their tool on our dataset.
Keshani et al. (Keshani et al., 2024) automates the process of reproducing Maven projects by generated buildspec files. They show that their approach simplifies generated buildspec files and also fixes issues in the currently existing buildspec files in Reproducible Central dataset. Our work also finds that a mitigation strategy to fix unreproducibility issue is by fixing the build script. In addition, we perform the first analysis of fixing by canonicalization.
Randrianaina et al. (Randrianaina et al., 2024) identifies the configuration options in that cause unreproducibility. The authors select highly configurable systems such as Linux, BusyBox, and ToyBox to study the unreproducibility of their builds. They find options for module signing, debug information, and profiling that cause unreproducibility. In our work, configuration options is one of the causes of unreproducibility. Especially, when the build process involves setting Maven profiles subsubsection 4.3.4.
Dietrich et al. (Dietrich et al., 2024b) show that only a 25% (out of 14,156) of the binaries in Maven ecosystem are actually bit by bit identical. These binaries come from 4 different registries - Maven, Google Assured Open Source Software, RedHat, and Oracle build-from-source. They propose 4 levels of binary equivalence to compare binaries which becomes increasingly lenient as we go from level 1 to level 4. First is to compare binaries bit by bit and this is the strictest level of binary equivalence check. Second is to compare binaries and ignore differences that have no semantic effect (for example, @Deprecated). OSS-Rebuild fits into the second level of checking binary equivalence. Third is to convert binaries into a common format and then compare them. jNorm fits into the third level of checking binary equivalence Finally, fourth is to compare binaries using a similarity metric. In the paper, they use tlsh (Oliver et al., 2013) to compare hash distance between two binaries; lesser the hash distance, more similar the binaries are. This work is similar to ours as they also evaluate effectiveness of jNorm over thousands of pairs of JVM bytecode (Dietrich et al., 2024a) to fix unreproducibility issues. However, we consider more causes of unreproducibility other than ones in JVM bytecode.
Finally, strip-nondeterminism  [^38] https://salsa.debian.org/reproducible-builds/strip-nondeterminism is a tool the Debian project that strips out timestamps and file system order from archives. However, it only takes care of these 2 causes of unreproducibility and does not cover the other causes of unreproducibility that we have discussed in this paper.
8. Conclusion
In this paper, we have presented the first large-scale study of build unreproducibility in Java. We have identified the six main causes of unreproducibility with underlying root causes: 1) build manifest, 2) software bill of materials, 3) filesystem, 4) JVM bytecode, 5) versioning properties, and 6) timestamps. Leveraging the same dataset, we have evaluated the effectiveness of canonicalization as a mitigation strategy to fix unreproducibility. Our results have shown that 26.89% of unreproducible artifacts can be canonicalized and become reproducible after canonicalization.
As future work, we want to formalize the notion of acceptable canonicalization by verifying what transformations can be applied to build artifacts without obscuring meaningful differences. This is important to ensure that the technique eliminates non-deterministic noise while preserving semantic integrity. Build reproducibility verification via canonicalization must be sound and precise.
9. Acknowledgements
We acknowledge the use of GitHub Copilot for assistance in polishing the writing of this paper. This work is supported by the CHAINS project funded by Swedish Foundation for Strategic Research (SSF), as well as by the Wallenberg Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
Appendix A Unreproducible MANIFEST.MF
The MANIFEST.MF file is a mandatory metadata file that contains information about the JAR file.
A.1. Built-By
This attribute is used to indicate the name of the user who built the JAR file. This can be different for environments as the user name is different.
Example Maven release org.apache.camel:came-ahc-ws:3.13.0 has an unreproducible artifact camel-ahc-ws-3.13.0.jar. This JAR file has a different Built-By attribute in the MANIFEST.MF file. In the reference version of the JAR file, the Built-By attribute is set to root while in the rebuild version of the JAR file, the Built-By attribute is set to aman.
Solution Remove the Built-By attribute from the MANIFEST.MF file.
A.2. Implementation-Build-Java-Vendor
This attribute records the vendor of the JDK used to build the JAR file. There are 17 different open source distributions of JDK  [^39] https://sdkman.io/jdks/ - GraalVM, OpenJDK, Temurin to name a few. Although all of them follow the same Java specification by Oracle, the implementation details could differ.
Example MANIFEST.MF of ldapchai-0.8.0.jar of Maven release com.github.ldapchai:ldapchai:0.8.0 shows difference in vendor when rebuilt. It changes from AdoptOpenJDK to OracleCorporation. AdoptOpenJDK, now known as Adoptium, indicates that the JAR file is built using Temurin. We use OpenJDK for the rebuild version, so the vendor is set to OracleCorporation as OpenJDK is maintained by Oracle.
Solution The buildspec file should document the exact vendor of the JDK.
A.3. Build-Jdk or Implementation-Build-Java-Version
This attribute reports the exact version of JDK used to build the JAR file. The JDK version given a fixed vendor is based on Semantic Versioning. Ensuring the exact same JDK version across different environments depends upon the major, minor, patch version and if this version has a pre-release flag or not. These four factors are in addition to the vendor of the JDK as discussed above.
Example Artifact camel-cloud-3.13.0.jar of Maven release org.apache.camel:camel-cloud:3.13.0 is built with different versions of JDK. The reference version of the JAR file has the Build-Jdk attribute set to 1.8.0\_292 while the rebuild version of the JAR file has the Build-Jdk attribute set to 1.8.0\_422. The difference in the JDK version is due to the update number (Clark and Reinhold, 2014).
Solution The buildspec file should document the exact version of JDK and provide an automated way to rebuild. There are some cases where the JDK version is precisely mentioned in Reproducible Central dataset. However, we remove them in our dataset filtering process as they require manual intervention to build.
A.4. Signed JARs
Java provides a way to sign each file in a JAR to ensure its users that the file has not been tampered with  [^40] https://docs.oracle.com/javase/tutorial/deployment/jar/intro.html.
Example The JAR file xmlchai-0.1.0.jar of Maven release org.jrivard.xmlchai:xmlchai:0.1.0 shows that signatures are removed in the rebuild version of the JAR file.
Solution Strip the signature from the JAR file as signatures can not be replicated.
A.5. Os-Version
This attribute is used to indicate the operating system version used to build the JAR file.
Example The artifact jandex-maven-plugin-3.1.0.jar from Maven release io.smallrye:jandex-maven-plugin:3.1.0 has different Os-Version attribute in the MANIFEST.MF file. It seems that the kernel version is different as the value changes from 5.15.0-1035-azure to 5.15.0-118-generic.
Solution This attribute should be stripped from the MANIFEST.MF file before comparison as it is environment specific and Reproducible Builds are environment independent.
A.6. Created-By
These attribute is used to indicate the name and version of the build tool or the Java version used to produce the JAR file. This can be created by Java jar tool or any build tool like Maven, Gradle, or Ant.
Example jline-console-3.22.0-sources.jar of Maven release org.jline:jline-console:3.22.0 has different Created-By attribute in the MANIFEST.MF file. The reference version reports ApacheMaven while the rebuild version reports MavenSourcePlugin3.2.1.
Solution The buildspec file should document the exact variant and version of the build tool used to build the JAR file as this ensures the same build tool is used across different environments.
A.7. Originally-Created-By
This attribute is added by maven-bundle-plugin when the JAR file is created again without cleaning the build directory. The plugin updates the JAR with this attribute to indicate the original build plugin that created the JAR file.
Example Maven release ch.qos.logback:logback-access:1.3.0-alpha14 has an artifact logback-access-1.3.0-alpha14.jar which does not have the Originally-Created-By when rebuilt.
Solution The buildspec file should document prerequisites to build the JAR file for the same reasons as Created-By attribute.
A.8. SCM-Revision
MANIFEST file can contain information related to version control system like Git. SCM-Revision attribute is used to indicate the commit hash of the source code used to build the JAR file  [^41] https://www.mojohaus.org/buildnumber-maven-plugin/usage.html.
Example The JAR file io.wcm.tooling.commons.crx-packmgr-helper-2.0.2.jar of Maven release io.wcm.tooling.commons:io.wcm.tooling.commons.crx-packmgr-helper:2.0.2 has different SCM-Revision attribute in the MANIFEST.MF file. The value changed from c48c286dat2022-01-06T11:46:12+01:00 to bc9cc174at2024-10-17T02:14:22Z. There are two reasons of differences here. First, the commit hash is different which indicates that commit hash corresponding to the git tag has changed. Second, the timestamp is different and this is due to embedding build time in the same attribute.
Solution The commit hash should remain the same across different builds as it guarantees that the source code is the same across different builds. buildspec files currently relies on Git tags to fetch source code, but they are mutable and can be moved to different commits. Thus, the buildspec file should document the exact commit hash of the source code.
A.9. SCM-Git-Branch
This attribute is used to indicate the branch of the source code used to build the JAR file.
Example MANIFEST.MF of ldapchai-0.8.0.jar of Maven release com.github.ldapchai:ldapchai:0.8.0 shows that the value changes from master to 338023a44e4dc62aff8985ca42c2f6743258b1c0.
Solution Branches are mutable in Git so they should not be embedded in the MANIFEST.MF file. Commit hash should be used instead as they are fixed.
A.10. Bnd-LastModified
This attribute is part of the OSGi specification to create bundles.
Example shiro-cas-1.9.0.jar is an artifact of org.apache.shiro:shiro-cas:1.9.0 where the value of this attribute changes from 1647431421514 to 1735860648961.
Solution Apache Maven provides a property \${project.build.outputTimestamp} that fixes the timestamp and can be read from again in subsequent rebuilds  [^42] https://maven.apache.org/guides/mini/guide-reproducible-builds.html. This property should can either be reused to set the value of Bnd-LastModified attribute or the attribute should be stripped from the MANIFEST.MF file.
A.11. Order of values for attributes
If the value of an attribute is a comma separated list, the order of the values can differ across different builds.
Example slf4j-api-2.0.6.jar of Maven release org.slf4j:slf4j-api:2.0.6 has an attribute Export-Package in the MANIFEST.MF file. This attribute is part of OSGi specification and is used to export packages that can be imported by other bundles. The value of this attribute is a list of packages whose order differs across the different builds.
Solution The attributes where this problem exists are Include-Resource, Private-Package, and Provide-Capability. The solution to depend upon the latest version of maven-bundle-plugin as it makes the order deterministic by sorting the values  [^43] https://github.com/apache/felix-dev/commit/d885d99a6a16660f655a4fd18e8a1a39beef0a15.
A.12. Addition or removal of attributes
Presence or absence of an attribute in the MANIFEST.MF file can also cause unreproducibility.
Example apache-any23-csvutils-2.7-sources.jar of Maven release org.apache.any23:apache-any23-csvutils:2.7 removed 10 attributes from the MANIFEST.MF file. These attributes are Specification-Title, Specification-Version, Specification-Vendor, Implementation-Title, Implementation-Version, Implementation-Vendor, Implementation-Build, Implementation-Build-Date, X-Compile-Source-JDK, and X-Compile-Target-JDK.
Solution We suggest to canonicalize these attributes. We cannot confirm what build tool embeds these attributes.
Appendix B Unreproducible pom.properties
The pom.properties file is an automatically generated file that contains information about the Maven project.
B.1. Order of properties
The file contains properties groupId, artifactId, version. However, their order can differ across different builds.
Example The pom.properties file of Maven release io.dropwizard.metrics:metrics-annotation:4.1.20 has the property version in different order  [^44] https://github.com/chains-project/reproducible-central/issues/17#issuecomment-2576045821.
Solution The order of the properties should be fixed in the pom.properties file or the file should be removed from the JAR file as it is not required for the JAR file to run.
B.2. Timestamp
While generating pom.properties file, Maven uses the current time to set the time of generation.
Example ch.qos.logback.db:logback-classic-db:1.2.11.1 embeds timestamp in one of the comments in the pom.properties file.
1 #Generated by Apache Maven
2 #Wed Apr 20 20:27:33 CEST 2022
3 #Wed Jan 29 08:18:40 UTC 2025
Solution The timestamp should either be set to the value of SOURCE\_DATE\_EPOCH or removed from the pom.properties file.
B.3. Eclipse properties
If a Maven module is built using, m2e - Maven build tool for Eclipse IDE, it adds some properties to the pom.properties file other than the ones added by standalone Maven build tool.
Example Maven release org.spdx:spdx-maven-plugin:0.7.0 has m2e.projectLocation and m2e.projectName properties in the pom.properties file.
Solution These attributes can be removed from the pom.properties file as they are not required for the JAR file to run.
Appendix C Unreproducible Software Bill of Materials
C.1. Removal of hash algorithms
CycloneDX supports multiple hash algorithms to calculate the hash of components. Our dataset shows that some rebuilds omits some of the hash algorithms.
Example ratis-3.1.1-cyclonedx.json artifact of Maven release org.apache.ratis:ratis:3.1.1 has different hash algorithms in the CycloneDX SBOM. The rebuild version of the SBOM omits the SHA3-384, SHA3-256, and SHA3-512 hash algorithms.
Solution These hash algorithms are not present in JDK 1.8.0_422 which is used to rebuild the artifact in our dataset. The solution is to use a JDK that has backported these hash algorithms. This has already been solved by modifying the buildspec file to use Azul JDK  [^45] https://github.com/jvm-repo-rebuild/reproducible-central/commit/ff0cc3037b7a1d0f009d4e4461c10f817c8658c7.
C.2. Addition or removal of components
SBOM records the components used to build the software. In the context of Maven, this can either be a dependency or a Maven submodule. Difference in components in SBOM mean that the dependencies are different across different builds and hence the build is unreproducible.
Example Maven release net.sourceforge.pmd:pmd:7.0.0 has an artifact pmd-7.0.0-cyclonedx.json which shows that the components attribute is different across reference and rebuild versions of the SBOM. The rebuild version documents two more Maven submodules which are net.sourceforge.pmd:pmd-cli:7.0.0 and net.sourceforge.pmd:pmd-dist:7.0.0.
Solution The build script here  [^46] https://github.com/pmd/pmd/blob/main/.ci/build.sh#L81 shows ad-hoc configuration of the build process. This should not be done as it hinders automation when checking for reproducibility. The fix here is to stick to conventional builds.
Example Maven release io.dropwizard.metrics:metrics-core:4.1.32 has a CycloneDX SBOM, metrics-core-4.1.32-cyclonedx.json, where all components are deleted except org.slf4j:slf4j-api:1.7.36.
Solution It is unclear why all components are deleted. The original dataset does not have the same result  [^47] https://github.com/jvm-repo-rebuild/reproducible-central/blob/55c06c8c4a080e66267f573560166c42071b2814/content/io/dropwizard/metrics/metrics-parent-4.1.32.diffoscope asked here:https://github.com/dropwizard/metrics/issues/4702. The solution that Reproducible Central uses is to ignore checking for generated SBOMs.
C.3. Modification of components
Even if the components are the same across different builds, the content of the components should exactly match across different builds.
Example Maven release org.apache.bcel:bcel:6.8.1 has an artifact bcel-6.8.1-cyclonedx.json where one of the components org.apache.commons:commons-lang3:3.14.0 has different hashes. SHA1 hash in reference versions is 29a8e03 while in the rebuild version is 1ed4711. The rebuild version’s hash is also consistent with the one on Maven Central  [^48] https://repo1.maven.org/maven2/org/apache/commons/commons-lang3/3.14.0/commons-lang3-3.14.0.jar.sha1. This happens because the local  .m2 folder has locally installed version which is referenced in the SBOM. However, the rebuild version uses the version from Maven Central.
Solution The release command should clean the local Maven repository before building the artifact. This can be done using purge-local-repository goal of maven-dependency-plugin  [^49] https://maven.apache.org/plugins/maven-dependency-plugin/examples/purging-local-repository.html.
C.4. metadata.timestamp
An attribute in the CycloneDX SBOM is used to indicate the time when the SBOM is generated. This can be different across different builds as the build time is different.
Example The CycloneDX SBOM, cyclonedx-core-java-7.3.2-cyclonedx.json of Maven release org.cyclonedx:cyclonedx-core-java:7.3.2 has different metadata.timestamp attribute.
Solution It should respect the Maven property \${project.build.outputTimestamp} to set the value of metadata.timestamp attribute. Although this has been addressed in pull request  [^50] https://github.com/CycloneDX/cyclonedx-core-java/pull/63 in release 3.0.8, but it does not seem to resolve the issue until release 8.0.0. Perhaps, a better way to deal with this is to canonicalize the timestamp attribute.
C.5. serialNumber
A unique serial number is assigned to an SBOM even if the content of the SBOM is the same  [^51] https://cyclonedx.org/docs/1.6/json/#serialNumber.
Example Maven release org.cyclonedx:cyclonedx-maven-plugin:2.7.5 produces cyclonedx-maven-plugin-2.7.5-cyclonedx.xml where difference in serialNumber attribute is observed.
Solution The serialNumber attribute should be removed before comparison as it is not relevant to the content of the SBOM. Else, the generation should depend upon some fixed content of the SBOM to ensure the same serial number across different builds. This has been addressed by computing serialNumber on the basis of groupId, artifactId, and version  [^52] https://github.com/CycloneDX/cyclonedx-maven-plugin/pull/425.
C.6. System Package Data Exchange (SPDX)
SPDX is a standard to document the licenses of the components used to build the software. It has non-deterministic attributes like created and licenseListVersion which can cause unreproducibility. created attribute is used to indicate the time when the SPDX file is created. licenseListVersion attribute is used to indicate the version of the SPDX license list used to document the licenses of the components. This list is updated externally and can vary with time.
Example We only see one instance of unreproducibility that falls under this category. Maven release org.apache.commons:commons-parent:60 has an artifact commons-parent-60.spdx.json where value of created and licenseListVersion attribute is different across different builds.
Solution These attributes should be removed before comparison as they are non-deterministic.
Appendix D Unreproducible Filesystem
D.1. Files are removed or added
If archives have different number of files, then the archive cannot be reproducible as the content of the archive is different.
Example Maven release commons-daemon:commons-daemon:1.4.0 has an unreproducible artifact commons-daemon-1.4.0-bin-windows.zip which does not contain 3 Windows executable binary in the rebuild version of the archive. There is custom non-Maven command that has to be executed during the build process to generate and place these binaries under the build folder. Since the command is not documented in the buildspec file, the rebuild version of the archive does not contain these binaries.
Solution The build process should be fixed to incorporate generation of other binaries along with the Maven build command. If this is not possible, the rebuilder should document the steps for generating the binaries in the buildspec file.
Example Maven release org.apache.rat:apache-rat-project:0.16.1 has an unreproducible artifact apache-rat-0.16.1-src.tar.bz2 which has an empty directory apache-rat-0.16.1/apache-rat-api/ compressed in the archive. This directory has been deleted in this version [^53] https://github.com/apache/creadur-rat/blob/c5a31fecc3b6d3697e20cb867e82c55decf969be/RELEASE-NOTES.txt#L22. However, it existed in the earlier versions and while checking out only the files are deleted but directory is kept as git does not track directories.
Solution The rebuilder should remove the empty directories from the source code before building the artifact. gitclean-df can be used to remove the empty directories from the source code.
Example Maven release org.apache.activemq:apache-artemis:2.28.0 has source archive, apache-artemis-2.28.0-source-release.tar.gz, as its artifact. It has pushed the node modules in the archive which are not present in the rebuild version.
Solution The builder should avoid pushing the node modules in the source archive and should only push the files that are tracked by the version control system. Ideally, the node modules are always untracked to reduce the size of git repository.
D.2. File permissions
The file in the archive can be created with different permissions across different builds. umask utility controls the permission of newly created files and directories and it can be different across different environments.
Example Maven release io.github.albertus82:unexepack:0.2.1 has an unreproducible artifact unexepack-0.2.1-bin.zip which has different file permissions for the file unexepack-0.2.1/unexepack.jar in the rebuild version of the archive.
Solution The rebuilder can either set the file permissions to a fixed value or set umask in the rebuilder script.
D.3. File ownership
The file in the archive can be created with different ownership across different builds.
Example Maven release org.apache.accumulo:accumulo:1.10.2 has an artifact accumulo-1.10.2-src.tar.gz. The file ownership has changed from christopher to aman.
Solution File ownership should be set to a fixed value by the rebuilder from all artifacts before comparison.
D.4. File timestamps
Example org.apache.zookeeper:parent:3.8.1 upon rebuilding produces parent-3.8.1.tar.gz where the timestamp of the files in the tar ball is set to the time of rebuild.
Solution The timestamp can either be set to a fixed value or stripped from the file by rebuilder before comparison.
D.5. File type
The file type of archive are different across different builds. This causes the size of the archives to be different as different compression algorithms could be used.
Example Consider Maven release io.github.git-commit-id:git-commit-id-maven-plugin:6.0.0 which has an unreproducible artifact git-commit-id-maven-plugin-6.0.0-sources.jar. It shows that the type of the file changes from Ziparchivedata,atleastv1.0toextract,compressionmethod=store to Javaarchivedata(JAR) [^54] Seems different from RC https://github.com/jvm-repo-rebuild/reproducible-central/blob/329fd7e4d7721cf7ea85fe203fadfb884be13fa9/content/io/github/git-commit-id/git-commit-id-maven-plugin-6.0.0.diffoscope#L129-L130. This seems to happen because outdated plugins are used to generated the source archive as it has attributes like Built-By in the MANIFEST.MF file which has been removed in the newer version of maven-archiver plugin  [^55] https://issues.apache.org/jira/browse/MSHARED-799.
Solution The builder should always use the specified version of the plugins to generate the archives. To help with this, the local repository can be cleaned before building the artifact using purge-local-repository  [^56] https://maven.apache.org/plugins/maven-dependency-plugin/examples/purging-local-repository.html.
Example
D.6. File size
Example org.apache.maven:apache-maven:3.8.1 has the executable mvn whose size is different across different builds. It changes from 5741 bytes to 5940 bytes because of different new lines in both executables.
Solution The rebuilder should replace the new lines from all files to either LF or CRLF before comparison based on the operating system of the builder. Builder should not replace the new lines as it could break behavior of the executable on different platforms  [^57] https://cwiki.apache.org/confluence/pages/viewpage.action?pageId=74682318.
D.7. File names
Example org.apache.jena:jena-permissions:4.3.2 has an unreproducible artifact jena-permissions-4.3.2-sources.jar where the file name is different in the rebuild version of the archive, but the content is identical.
Solution The builder should not embed system specific paths in any artifact (Andersson, 2024). Go ecosystem has a tool trimpath which removes the system specific paths.
D.8. File paths embedded
If absolute file paths are embedded in the archive, then the build is unreproducible. In our dataset, we observe that the absolute file paths are embedded in JVM bytecode and configuration files.
Example Maven release org.apache.cxf.fediz:common:1.6.1 has an artifact common-1.6.1.jar which has a file path embedded in the JVM bytecode.
Example Maven release org.owasp:dependency-check-cli:7.0.4 embeds absolute paths in dependencycheck-cache.properties.
Solution The builder should not embed absolute paths in the configuration files and stick to relative paths.
Appendix E Unreproducible JVM bytecode
E.1. Debug information
This corresponds to the information in the JVM bytecode that is used for debugging the Java program. There are three types of debug information in the JVM bytecode - SourceFile, LineNumberTable, and LocalVariableTable.
Example In release io.dropwizard.metrics:metrics-httpclient:4.1.27, the artifact metrics-httpclient-4.1.27.jar shows an extra line in LineNumberTable which corresponds to the return JVM bytecode instruction for a void method. This could be because the precise version and vendor of JDK is not used.
Solution The rebuilder should remove the information related to debugging from the JVM bytecode before comparison.
E.2. Optimizations or de-optimizations
The JVM bytecode can be optimized or de-optimized across different builds if the JDK version used to build the artifact is different. Newer versions of JDK can have better optimizations while older versions can have less performant code. We have found the following types of differences in the JVM bytecode across different builds:
  • try-with-resource is simplified with less goto instructions and close() invocations. Example: io.dropwizard.metrics:metrics-servlets:4.1.20.
  • return type in invokevirtual change from java/nio/Buffer to java/nio/ByteBuffer. Example: io.dropwizard.metrics:metrics-collectd:4.1.20.
  • invokeinterface is changed to invokevirtual. Example: com.github.ldapchai:ldapchai:0.8.6.
  • implementation finally is changed from using subroutines (jsr and ret) to using more inline code. Example: net.bytebuddy:byte-buddy-dep:1.14.5.
  • static and final modifiers are added or removed from anonymous classes. Example: dev.langchain4j:langchain4j-core:0.26.0.
  • invocations to Objects.requireNonNull are added in rebuild version of the JVM bytecode. Example: dev.langchain4j:langchain4j-core:0.26.0.
  • checkcast instructions are reduced. Example: org.apache.sling::org.apache.sling.servlets.resolver:2.8.2.
  • implementation of nested enums differ. This is due to a commit in OpenJDK 22  [^58] https://github.com/openjdk/jdk/pull/10797 which changes how switch map tables are generated.
  • synthetic method, values(), has bytecode under a separate method in reference version and under the static initializers in rebuild version  [^59] https://bugs.openjdk.org/browse/JDK-8241798. Example: com.google.guava:guava:32.0.0-jre
  • synthetic field for inner class is added in rebuild version  [^60] https://bugs.openjdk.org/browse/JDK-8271717. Example: com.google.guava:guava:32.0.0-jre
  • local variables in LocalVariableTable are in the scope for a more number of instructions. Example: org.apache.helix:helix-common:1.0.3.
  • ldc\_w instructions are replaced with ldc instructions. Example: org.codehaus.mojo:jaxb2-maven-plugin:3.1.0.
  • strings are concatenated which reduces the number of ldc instructions. Example: org.apache.sling.event:org.apache.sling.event:4.2.18.
  • StackMapTable is modified. org.apache.sling:org.apache.sling.servlets.resolver:2.8.2
Solution The rebuilder should use the precise version of JDK to build the artifact as the reference version. This is a non-trivial task as the exact version of JDK is not always mentioned and there are a myriad of JDK versions and vendors available. As an example, we found that building com.github.ldapchai:ldapchai:0.8.6 with JDK 22 solves all differences related to JVM bytecode.
E.3. Refactoring
These changes simply change the structure of the JVM bytecode without altering the semantics or performance. Changes in order of sections in JVM bytecode is an example of refactoring. The JVM bytecode contains sections that are order independent, but they can differ across different builds and fail the build reproducibility check. In our dataset, we observe that the order of methods, fields, static initializers, entries in the constant pool, array type values in annotation, method declaration in interfaces, and order of inner classes can differ across different builds.
Example In Maven release org.apache.activemq:activemq-runtime-config:5.16.4, the artifact activemq-runtime-config-5.16.4.jar uses jakarta.xml.bind.annotation.XmlElementRefs which takes an array of XmlElementRef annotations. The order of these values in the array is different across different builds.
Solution The rebuilder should sort the order independent sections of the JVM bytecode before comparison.
Apart from order, we observe 2 types of refactoring that did not change the order.
  • private modifier is removed from a synthetic field in an anonymous class in artifact antisamy-1.7.3.jar. We consider this as a refactoring as fields in anonymous classes are implicitly private.
  • The positive and negative branches of an if statement are swapped in the JVM bytecode of Maven release eu.maveniverse.maven.mima:2.4.2:standalone-static-uber
Solution The rebuilder should canonicalize the JVM bytecode before comparison.
E.4. Embedded data
Absolute file paths, timestamp, java version, project version and other environment related variables are embedded in the JVM bytecode.
Example Maven release org.apache.hive:hive-exec:4.0.0-alpha-2 embeds git branch name, user name, timestamp, and git repository path in hive-exec-4.0.0-alpha-2.jar which makes it unreproducible.
Solution The builder should embed such environment related variables in the JVM bytecode.
E.5. Build time generated/modified code
The Java compiler produces synthetic code during build in order to handle access to members (Rimenti, 2018). The compiler also adds names to lambda functions (Evans, 2020) and inner classes (Maverik, 2016) which are not present in the source code. In addition to these capabilities, Maven provides plugins to generate or modify code during the build process. All of this generated code can be different across different builds and hence the build is unreproducible.
Example Maven release io.github.chains-project:maven-lockfile-github-action:3.4.0 generates a lot of code during the build process. The code is generated by two dependencies of maven-lockfile, o.quarkiverse.githubaction:quarkus-github-action:2.0.1 and io.quarkus:quarkus-arc:3.1.1.Final in order to make the application compatible with GitHub actions. The generated code differs in order of fields, different constant values, and embedded system properties. Another example is org.apache.hive:kafka-handler:4.0.0-alpha-2 which uses maven-shade-plugin to rename package names from kafkaesqueesque to kafkaesque.
Example Maven release org.apache.helix:zookeeper-api:1.0.3 has an artifact zookeeper-api-1.0.3.jar. The lambda function changes from lambda\\$parseRoutingData\\$19 to lambda\\$parseRoutingData\\$1.
Example Maven release org.apache.hive:hive-parser:4.0.0-alpha-2 has an artifact hive-parser-4.0.0-alpha-2-sources.jar that has sources generated by ANTLR3 Maven plugin. The generated code has comments where order of values differs across builds. This issue is fixed by sorting the values in the comments  [^61] https://github.com/antlr/antlr4/pull/3809.
Example Maven release org.apache.shiro:shiro-aspectj:1.11.0 produces an artifact where AspectJ metadata is embedded in the JVM bytecode. The metadata does not exist in the reference build. This issue has been fixed by forcing the compilation of AspectJ classes  [^62] https://github.com/apache/shiro/commit/2d92460cd8b7a621be0b82725b7a5cd6952734c1.
Example Maven release io.fabric8:kubernetes-model
-apiextensions:6.4.0
contains build time generated code where the order of if-else blocks is different across different builds. We do not classify them under refactoring as the order of if-else blocks is relevant to the semantics of the code.
Solution The generated code should be deterministic across different builds. This calls for fixes in the Java compiler or the Maven plugins that generate code during the build process.
Appendix F Unreproducible Timestamps
F.1. Properties file
The properties files in a JAR file stores configuration related to the project. This can be either a file manually created by the developer or generated by the build tools. In our dataset, we observe that the timestamps in the properties files are different across different builds.
Example Maven release com.adobe.acs:acs-aem-commons-bundle:5.2.0 has files embedded in the JAR file that have different timestamps of generations across different builds.
Solution The plugin that generates the properties file should respect the Maven property \${project.build.outputTimestamp} to set the value of the timestamp in the properties file.
F.2. Documentation
The generated documentation can have timestamps embedded in it.
Example Maven release org.apache.synapse:synapse-documentation:3.0.2 has an artifact synapse-documentation-3.0.2.jar which has a timestamp embedded in the documentation. org.apache.karaf:manual:4.4.0 is another example.
Solution The plugin that generates the documentation should respect the Maven property \${project.build.outputTimestamp} to set the value of the timestamp in the documentation.
F.3. Shell scripts
Example Maven release org.apache.sling:org.apache.sling.feature.cpconverter:1.3.4 has an artifact org.apache.sling.feature.cpconverter-1.3.4.tar.gz that packs a shell script bin/cp2feature. The script takes build timestamp as one of the arguments and the timestamp is different across different builds.
Solution The shell script should respect the Maven property \${project.build.outputTimestamp} or the standard SOURCE\_DATE\_EPOCH environment variable to set the value of the timestamp in the shell script.
F.4. Binaries
Example Maven release io.github.albertus82:unexepack:0.2.1 embeds a Windows executable that embeds the timestamp.
Solution The build tool for binary should respect SOURCE\_DATE\_EPOCH environment variable to set the value of the timestamp in the binary.
F.5. SBOM
Unreproducibility due to attribute timestamp in the SBOM as explained in section C.4.
F.6. JVM bytecode
JVM bytecode can have timestamps embedded as values of variables.
Example package-info.class in 4.0.0-alpha-2:hive-standalone-metastore-server:4.0.0-alpha-2 has timestamp embedded in an annotation.
Solution This should respect the Maven property \${project.build.outputTimestamp} to set the value of the timestamp in the JVM bytecode.
F.7. File timestamps
Files have timestamp for creation or modification time as shown in section D.4.
F.8. MANIFEST
Attribute Bnd-LastModified tends to be non-deterministic as documented in section A.10.
F.9. NOTICE
Example Maven release org.apache:apache:24 has a sources Jar where the year of build is embedded in the apache-24/NOTICE file.
Solution The plugin that generates the NOTICE file should respect the Maven property \${project.build.outputTimestamp} to set the value of the timestamp in the NOTICE file.
F.10. Servlets
Example Maven release org.apache.nifi:nifi-web-ui:1.27.0 has embedded timestamps. These timestamps are created by Jasper which compiles JSP files to servlets.
Solution The plugin that generates the servlets should respect the Maven property \${project.build.outputTimestamp} to set the value of the timestamp in the servlets.
Appendix G Unreproducible Order
G.1. Ordering in JVM bytecode
The JVM bytecode contains sections that are order independent, but they can differ across different builds and fail the build reproducibility check. In our dataset, we observe that the order of methods, fields, static initializers, entries in the constant pool, array type values in annotation, method declaration in interfaces, and order of inner classes can differ across different builds.
Example In Maven release org.apache.activemq:activemq-runtime-config:5.16.4, the artifact activemq-runtime-config-5.16.4.jar uses jakarta.xml.bind.annotation.XmlElementRefs which takes an array of XmlElementRef annotations. The order of these values in the array is different across different builds.
Solution The rebuilder should sort the order independent sections of the JVM bytecode before comparison
G.2. pom.properties
The attributes in the pom.properties file are not in the same order across different builds.
Example io.dropwizard.metrics:metrics-jcache:4.1.20 has an artifact metrics-jcache-4.1.20.jar which has a version attribute that appears at line 1 in the reference version and at line 3 in the rebuild version. io.takari.maven:takari-smart-builder:1.0.0 is another example where order of artifactId is different.
Solution The rebuilder should sort the attributes in the pom.properties file before comparison. However, this file is most commonly generated by maven-archiver-plugin and this ordering issue has been fixed  [^63] https://github.com/apache/maven-archiver/commit/763a940540eefad74f9ba73cb5eed288dc4e639d.
G.3. XML files
Example Order of remoteResources in file generated by maven-remote-resources-plugin is different across different builds. This is seen in org.apache.axis2:axis2-resource-bundle:1.8.2.
Solution The rebuilder should sort the order of elements in the XML files before comparison. This won’t affect the future releases as this unstable order has been fixed in 3.3.0 version of maven-remote-resources-plugin  [^64] https://github.com/apache/maven-remote-resources-plugin/pull/74.
Example components.xml generated by plexus-containers has different order of elements across different builds.
Solution The rebuilder should sort the order of elements in the XML files before comparison. This has been fixed recent releases of plexus-containers  [^65] https://github.com/codehaus-plexus/plexus-containers/issues/8.
G.4. JMH benchmark
JMH is a microbenchmarking framework for Java.
Example org.apache.hive:hive-metastore-benchmarks:4.0.0-alpha-2 publishes output from JMH benchmarks in the JAR file. Even though the contents remain the same, the order of results is different across different builds.
Solution Builder should not publish the output from JMH benchmarks in the JAR file. It is a metric that is used to improve the performance of the code and should not be part of the released artifact.
G.5. Files in archive
Jar file is a zip file and the order of files can be in any order.
Example Order of files in org.apache.synapse:synapse-package-archetype:3.0.2 do not stay the same across builds.
Solution The order of zip files can be canonicalized by rebuilder before comparison. For future releases, this has been fixed  [^66] https://issues.apache.org/jira/browse/MJAR-263  [^67] https://issues.apache.org/jira/browse/MSHADE-347.
G.6. Generated files
Example Maven release org.apache.hive:hive-parser:4.0.0-alpha-2 generates a file using ANTLR3 Maven plugin which creates a different order of values in the comments across different builds.
Solution The plugin that generates the file should sort the order of values in the comments before comparison. ANTLR3 is not maintained anymore but this issue has been fixed in ANTLR4  [^68] https://github.com/antlr/antlr4/pull/3809.
Example net.sourceforge.pmd:pmd-cli:7.0.0-rc3 generates a shell completion script which has different order of flags across different builds.
Solution The plugin that generates the file should sort the order of flags in the shell completion script before comparison. This has been fixed in later version of pmd-cli.
Example Apache Tomcat JspC creates web.xml file with Java Server Pages configuration. We observe that the order declaration is different across builds in Maven release org.apache.hive:hive-service:4.0.0-alpha-2.
Solution The plugin that generates the file should sort the order of declarations in the web.xml file before comparison.
G.7. Attribute values in MANIFEST
Some values are comma separated lists whose order can differ as shown in section A.11.
Appendix H Mysteries
Appendix I Optimizations not canonicalized by jNorm
The first case is how implementation enums differs between JDK 17 and JDK 22 compiled artifacts. JDK 17 creates a lookup table to store the values of the enum while JDK 22 simply stores it based on the order of the enum values  [^72] https://github.com/openjdk/jdk/pull/10797. We observe this difference in com.github.ldapchai:ldapchai:0.8.6 as the reference version is built with JDK 22 and the rebuilt version is built with JDK 17.
The second case is the usage of invokevirtual in JDK 17 and invokeinterface in JDK 21 bytecode instructions to execute toString, equals, and getClass methods. They differ in the time the method is resolved. The invokeinterface instruction resolves the method at runtime while invokevirtual resolves the method at compile time. We observed this difference in our dataset in dev.langchain4j:langchain4j:0.28.0 when toString method is invoked on instance of class.
The third case that jNorm is unable to canonicalize the implementation of string concatenation when its operands require a cast to String. For example, "ErrordeterminingJDBCtypeforcolumn"+column+".Cause:"+e is a string concatenation expressions in artifact mybatis-3.5.16.jar. jNorm identifies that there are two ways to implement string concatenation (Shipilev, 2024) - using StringBuilder (JDK < 9) and using invokedynamic (JDK $\geq$ 9) instruction and canonicalizes the expression to the second one. However, it fails to canonicalize the string conversion of e which is an instance of java.sql.SQLException before concatenation using invokedynamic.
Implementation of try and try-with-resource statements and finally clauses also differ across Java versions and is also not canonicalized by jNorm. For example, reference and rebuild artifacts of io.dropwizard.metrics:4.1.20:metrics-servlets-4.1.20 are compiled with JDK 8 and JDK 11 respectively. The rebuild artifact is optimized in the way it handles closing of resource. Another example is net.bytebuddy:1.14.5:byte-buddy-1.14.5.jar that uses deprecated bytecode instructions jsr and ret to implement closing of resource.
jNorm is unable to handle differences in how outer class reference is handled in the bytecode. A synthetic field called this\\$0 is created in the bytecode to refer to the outer class from inner class. The reference version of com.google.guava:guava:32.0.1-jre is built using custom JDK which omits this synthetic field  [^73] https://bugs.openjdk.org/browse/JDK-8271717 as it is not needed. We also observe that patch versions of JDK 17 handle this differently. They refer to the outer class reference using either this.this\\$0 or this\\$0. The former adds another getfield instruction to the bytecode. psi-probe-tomcat85-3.7.0-tests.jar is an example of artifact that contains this difference.

### v2

@@
-Canonicalization for Unreproducible Builds in Java
+Causes and Canonicalization of Unreproducible Builds in Java
 Abstract.
@@
 We study actionable mitigations: artifact and bytecode canonicalization using OSS-Rebuild and jNorm respectively.
-Finally, we present Chains-Rebuild, a tool that raises reproducibility success from 9.48% to 26.89% on 12,283 unreproducible artifacts.
+Finally, we present Chains-Rebuild (improvements to OSS-Rebuild), a tool that raises reproducibility success from 9.48% to 26.60% on 12,803 unreproducible artifacts.
 To sum up, our contributions are the first large-scale taxonomy of build unreproducibility causes in Java, a publicly available dataset of unreproducible builds, and Chains-Rebuild, a canonicalization tool for mitigating unreproducible builds in Java.
@@
 The growing complexity of software supply chains (Soto-Valero et al., 2022; Cox, 2019), coupled with the increasing frequency of supply chain attacks  [^1] https://www.sonatype.com/hubfs/SSCR-2024/SSCR_2024-FINAL-10-10-24.pdf, raises concerns about software integrity.
-In such a fragmented ecosystem, relying solely on assumptions about the identity of the distributor (Thompson, 1984) is no longer sufficient - what is needed is verifiable evidence that the software one installs corresponds exactly to its declared source.
+In such a fragmented ecosystem, relying solely on assumptions about the identity of the distributor (Thompson, 1984) is no longer sufficient — what is needed is verifiable evidence that the software one installs corresponds exactly to its declared source.
 This challenge is especially acute in open source environments, where binaries are often distributed separately from their source code (Davies et al., 2011), making it difficult for users to independently validate what they are running.
@@
 Any backdoor or malicious code would be detected by the verifier as the built artifact would not match the original artifact.
-This approach has already seen adoption in security-critical environments, such as in certain federal government contracts, where vendors are required to supply source code so the agencies can perform their own builds and verify the integrity of the software  [^2] https://news.ycombinator.com/item?id=43492115.
+This approach has already seen adoption in security and privacy critical environments, such as Bitcoin and Tor Browser (Levsen et al., 2025).
+Ensuring that clients for both of them are free of any backdoors injected by the build system.
 Despite the promises of Reproducible Builds, ensuring and verifying reproducibility at scale remains technically challenging due to the multitude of spurious differences in build outputs (e.g., timestamps, file ordering).
 Those differences make binaries appear different even when built from the same source, with an untampered build pipeline (Fourné et al., 2023).
-Spurious differences hinder reproducibility and make it difficult to justify the differences between two builds.
-XZ Utils is a widely used compression library that is used in almost all Linux distributions.
-In 2024, it is found that the XZ Utils has a backdoor to give remote code execution to the attacker (Przymus and Durieux, 2025).
-However, this backdoor is only observed in the official tarballs on the registry and do not exist in the git repository.
-Such supply chains attacks can be prevented if build process is reproducible.
-In the case of XZ Utils, if the tarball on the package registry does not match the one built from the source code, then the user can be sure that the tarball is tampered with (Malka, 2025).
+Spurious differences hinder reproducibility and create a new attack surface.
+Several high-profile supply chain attacks have exploited unreproducible builds.
+For instance, in 2024 it was discovered that the official tarballs of XZ Utils, compression library used in almost all Linux distributions, contained a backdoor that enabled remote code execution, even though the backdoor was absent in the project’s Git repository (Przymus and Durieux, 2025).
+Similarly, the Ultralytics Python package, a widely used library for AI models, was compromised when attackers injected malicious code into the build pipeline, resulting in a PyPI release with a backdoor (Larson, 2024).
+A related incident targeted the Solana ecosystem as well (Claburn, 2024).
+These incidents share a common vector: the artifacts distributed through registries or release tarballs differed from what could be obtained by building directly from the publicly available source code.
+If the build processes for these projects had been reproducible, users or third party rebuilders could have verified that the distributed binaries and packages matched the source.
+Any discrepancies would have immediately signaled tampering, thereby preventing such attacks (Malka, 2025).
 In this work, we contribute to solving the problem of achieving build reproducibility in the context of enterprise software in Java.
@@
 We aim at building the most comprehensive taxonomy of unreproducible builds in Java, and the corresponding mitigation strategies.
-Furthermore, we perform a deep study how canonicalization - removal of non-deterministic differences - is a good solution for mitigating unreproducibility.
+Furthermore, we perform a deep study how canonicalization — removal of non-deterministic differences — is a good solution for mitigating unreproducibility.
 Our approach for analyzing unreproducible builds in Java is shown in Figure 1.
@@
 Together, these two levels of canonicalization address both semantic and non-semantic sources of non-determinism, significantly improving overall build reproducibility.
-Our study reveals 6 original categories of unreproducibility causes in Java, each with the underlying root causes and the best practice mitigations.
+Our study reveals 6 reasons for unreproducibility in Java, each with the underlying root causes and the best practice mitigations.
 For example, changes in the bytecode in Java artifacts.
 Our large-scale experiments empirically measure the effectiveness of two state-of-the-art canonicalization techniques.
-First, we show that bytecode canonicalization with jNorm (Schott et al., 2024) can successfully canonicalize 267 out of 898 unreproducible artifacts with JVM bytecode.
+First, we show that bytecode canonicalization with jNorm (Schott et al., 2024) can successfully canonicalize 275 out of 936 (29.38%) unreproducible artifacts with JVM bytecode.
 We study the reasons why jNorm fails to canonicalize the remaining artifacts, which are promising areas of future work.
-Next, we evaluate artifact canonicalization using Google’s OSS-Rebuild and find that it can successfully canonicalize 9.48% (1,165 / 12,283) of unreproducible artifacts.
-By implementing new canonicalization patterns informed by our taxonomy, our new tool Chains-Rebuild increases the verification success rate to 26.89% (3,303 / 12,283).
+Next, we evaluate artifact canonicalization using Google’s OSS-Rebuild and find that it can successfully canonicalize 9.41% (1,205 / 12,803) of unreproducible artifacts.
+By implementing new canonicalization patterns informed by our taxonomy, our new tool Chains-Rebuild increases the verification success rate to 26.60% (3,405/ 12,803).
 Our paper is the first to demonstrate, at a large scale, that canonicalization is a promising mitigation strategy for fixing build unreproducibility in the Java ecosystem.
@@
   • A novel, comprehensive taxonomy of 6 unreproducibility causes and their mitigation in Java, helping practitioners to improve build reproducibility in their projects.
-  • A large-scale experiment on the effectiveness of canonicalization, showing that it is a promising solution to mitigate unreproducibility in Java.
-Artifact canonicalization is more effective at scale than bytecode canonicalization.
-  • A publicly available, curated dataset of 7,961 unreproducible Maven releases for future research.
+  • A large-scale experiment on the effectiveness of canonicalization, showing that it is a promising solution to mitigate unreproducibility in Java, with artifact canonicalization being more effective at scale than bytecode canonicalization.
+  • A publicly available, curated dataset of 8,292 unreproducible Maven releases for future research.
 2. Framework for Build Reproducibility
@@
 We work in the context of the Java programming language and its Maven build system.
+In the following, we illustrate the key concepts with the example of Apache Guava.
 Package Registry: A package registry is a centralized repository of software packages that can be reused by other applications.
@@
 In Maven, a module is referenced by a combination of Group Identifier, Artifact Identifier, and Version.
-Group ID is a unique identifier for the organization, eg.
-“org.apache”   [^6] https://maven.apache.org/guides/mini/guide-reproducible-builds.html.
+Group ID is a unique identifier for the organization, e.g., “org.apache”   [^5] https://maven.apache.org/guides/mini/guide-reproducible-builds.html.
 Artifact ID is the unique name of the module and is unique within the group ID.
 For example, com.google.guava:guava:32.0.0-jre is a Maven module where com.google.guava is the group ID, guava is the artifact ID, and 32.0.0-jre is the version.
-There are two types of modules: a source module is a module that is hosted on a version control system (eg.
-Github), not necessarily published on a package registry; a released module is one published on a package registry, Maven Central in this paper.
-Both released and source modules can be identified by the combination of group ID, artifact ID, and version to refer to both types of module.
+There are two types of modules: a source module is a module that is hosted on a version control system (e.g., Github), not necessarily published on a package registry; a released module is one published on a package registry, Maven Central in this paper.
+Both released and source modules can be identified by the combination of group ID, artifact ID, and version.
 In the latter case, we assume the presence of a mapping from version to Git tag or Git commit.
@@
 It contains build configuration for all the modules in a so-called ‘POM file’.
-For example, Maven project guava contains 6 source modules - guava-parent, guava-tests, guava-testlib, guava-gwt, guava-bom, and guava.
+For example, the Maven project guava contains 6 source modules — guava-parent, guava-tests, guava-testlib, guava-gwt, guava-bom, and guava.
 Artifact: An artifact is any file that is part of the released module.
@@
 It also allows users to verify that the executable provided to them is free of backdoors as they can build the software themselves and compare the output with the provided executable.
-(1) $ReferenceArtifact\leftarrow Build(SourceModule)$
-(2) $RebuildArtifact\leftarrow Rebuild(SourceModule)$
-(3) $ReferenceArtifact=RebuildArtifact$
-There are 2 other terms in the literature that are used interchangeably with build reproducibility - “verifiable builds” and “accountable builds” (Pohl et al., 2025).
+[…]
+[…]
+[…]
+There are two other terms in the literature that are used interchangeably with build reproducibility — “verifiable builds” and “accountable builds” (Pohl et al., 2025).
 Builds are verifiable if the build process can be modified in order to make the output artifact bit-by-bit identical.
@@
 2.3. Mitigation
-We claim that any cause of unreproducibility can be mitigated in three ways - fixing the build process, fixing the rebuild process, or canonicalizing the output of the rebuild process.
+Our background study on fixing reproducibility (see  Subsection 8.3), shows that any cause of unreproducibility can be mitigated in three ways — fixing the build process, fixing the rebuild process, or canonicalizing the output of the rebuild process.
 Fixing the build process and Fixing the rebuild process involve modifying inputs such as source code, dependencies, or build tools (e.g., build scripts and configurations) to the build or rebuild workflows in order to eliminate sources of non-determinism.
@@
 3. Dataset of Unreproducible Builds
+Dataset Stage Projects Releases
+Count Count
+Initial Reproducible Central (d280bf1) 706 4,956
+After filters in Subsubsection 3.2.1 684 4,030
+After filters in Subsubsection 3.2.2 271 1,144
+Table 1. Evolution of dataset size through filtering stages, showing the reduction from initial Reproducible Central data to final unreproducible releases.
 3.1. The Reproducible Central Project
-The Reproducible Central project (Boutemy, 2024) is a rebuild infrastructure to verify whether Maven projects are reproducible.
+The Reproducible Central project (Boutemy, 2024) is a rebuild infrastructure to verify whether Java projects are reproducible.
 It specifies list of steps to build a Java project from source in a buildspec file  [^6] https://github.com/jvm-repo-rebuild/reproducible-central/blob/master/doc/BUILDSPEC.md.
@@
 We take a snapshot of the Reproducible Central dataset on 8th October, 2024 whose commit is d280bf1.
-At this date, Reproducible Central contains 706 Maven projects, their corresponding build scripts, and a history of 4,956 releases.
+At this date, Reproducible Central contains 706 projects, their corresponding build scripts, and a history of 4,956 releases.
 It is the largest dataset attempting rebuilds of Java projects and is also maintained by maintainers of Apache Maven.
-Our goal is to curate a dataset of unreproducible released Maven modules in order to study the reasons why unreproducible builds occur.
-In order to curate the dataset of unreproducible released modules, we first apply 2 filters before running the rebuilds and 1 filter afterwards.
+Our goal is to curate a dataset of unreproducible released Java modules in order to study the reasons why unreproducible builds occur.
+In order to curate the dataset of unreproducible released modules, we first apply one filter before running the rebuilds and two filter afterwards.
 3.2. Filtering
-3.2.1. Maven Projects
-First, we filter out Maven projects that require manual intervention to build so that we can run the entire rebuild process automatically.
+Table 1 shows the evolution of the dataset size through filtering stages.
+We start with 706 Java projects and 4,956 releases in the Reproducible Central dataset.
+It contains 4,851, 103, and 2 releases built with Maven [^7] https://maven.apache.org/, Gradle [^8] https://gradle.org/, and SBT  [^9] https://www.scala-sbt.org/, respectively.
+Next, we apply two filters to the dataset which reduces the dataset to 684 Java projects and 4,030 releases.
+Finally, we only keep the unreproducible releases and projects which results in 271 Java projects and 1,144 releases.
+Next, we account for the fact that the 271 Java projects are made of multiple modules.
+Consequently, we further split the 271 Java projects and their 1,144 releases into 8,292 unreproducible module releases as we account for unreproducibility per <GroupId, ArtifactId, Version> coordinates of Java modules rather than the entire source project.
+3.2.1. Filter Java Projects
+We filter out Java projects that require manual intervention to build.
+This is to ensure that we can run the entire rebuild process automatically.
 This is indicated in buildspec file either by the presence of build tool or when the rebuild command is prefixed with ‘SHELL’.
@@
 We exclude projects that do not have a POM file at the root directory as it also requires manual intervention to get the path to rebuild artifacts in build directory.
-Second, we only include projects that use Maven as their build tool (some projects use Gradle or SBT).
-This is because Maven is the most mature build tool for Java projects, most widely used in the Java ecosystem (Sulír and Porubän, 2016).
-Moreover, Maven builds success rate is higher than Gradle and Ant (Maes-Bermejo et al., 2022).
-Once all rebuilds are done, we exclude projects where the unreproducible artifacts are not a subset of the all artifacts released on Maven Central.
+Once all rebuilds are done, we apply the second filter that excludes projects where the unreproducible artifacts are not a subset of all the artifacts released on Maven Central.
 This can happen when not all artifacts produced by the build process are released on Maven Central.
 For example, content/org.infinispan:protostream-aggregator:5.0.7.Final produces an artifact rotostream-integrationtests-5.0.7.Final-test-sources.jar which is not released on Maven Central.
-After applying the filters, we end up with 3,927 releases over 665 Maven projects.
-3.2.2. Released Modules
+After applying the two filters, we end up with 4,030 releases over 684 Java projects.
+3.2.2. Filter Released Modules
 Next, we need to identify the released modules to compare against.
-In order to map each unreproducible project to corresponding released modules, we first identify the Maven modules in each project.
+In order to map each unreproducible project to corresponding released modules, we first identify the Java submodules in each project.
 org.apache.aries.cdi:org.apache.aries.cdi.executable:1.1.5 is an example of a source module that does not have a corresponding URL on Maven Central.
 We parse the POM file located at the root directory of the project as it contains the required information about the modules.
-For example, jpmml-python is a Maven project hosted on GitHub  [^8] https://github.com/jpmml/jpmml-python/tree/1.2.2 which has 3 Maven modules with version 1.2.2 - org.jpmml:jpmml-python, org.jpmml:jpmml-python:1.2.2, and org.jpmml:pmml-python.
-Thus, jpmml-python Maven project on Reproducible Central corresponds to 3 Maven modules in our dataset.
-2 out of 3 released modules in this project contain unreproducible artifacts.
+For example, jpmml-python is a Maven project hosted on GitHub  [^10] https://github.com/jpmml/jpmml-python/tree/1.2.2 which has three Maven modules with version 1.2.2org.jpmml:jpmml-python, org.jpmml:pmml-python:1.2.2, and org.jpmml:pmml-python-testing.
+Thus, jpmml-python Maven project on Reproducible Central corresponds to three Maven modules in our dataset.
+Two out of three released modules in this project contain unreproducible artifacts.
 org.jpmml:pmml-python-testing:1.2.2 has pmml-python-testing-1.2.2-sources.jar and org.jpmml:pmml-python:1.2.2 has pmml-python-1.2.2-sources.jar.
-Finally, there are 4 cases where unreproducible reference and rebuild artifact are not mapped correctly.
-Maven module org.apache.ratis:ratis-assembly:3.1.0 and 3 more of its versions have an artifact with .pom extension.
+Finally, there are four 4 cases where unreproducible reference and rebuild artifact are not mapped correctly.
+Maven module org.apache.ratis:ratis-assembly:3.1.0 and three more of its versions have an artifact with .pom extension.
 The rebuild script maps this artifact to the Jar file produced during rebuild.
-This is a bug in the Reproducible Central rebuild script and have reported this to the author  [^9] https://github.com/jvm-repo-rebuild/reproducible-central/issues/2164.
-At the end of the pipeline, we end up with 7,961 Maven modules which have a total of 34,271 artifacts. Out of 34,271 artifacts, there are 12,283 unreproducible artifacts (35.8%).
+The third and final filter excludes these four cases.
+This is a bug in the Reproducible Central rebuild script and we have reported this to the author  [^11] https://github.com/jvm-repo-rebuild/reproducible-central/issues/2164.
+At the end of the pipeline, we end up with 8,292 released Java modules which have a total of 35,956 artifacts from 4,030 releases of 684 projects.
+Finally, we end up with 271 Java projects and 1,144 releases which are unreproducible and they correspond to 8,292 unreproducible released modules that have 12,803/35,956 (35.8%) unreproducible artifacts.
 3.3. Description of the Dataset
-The dataset contains groupId, artifactId, version, and the URL of all 7,961 unreproducible released modules on Maven Central.
-It also contains the reference and rebuild artifact pair of all 12,283 unreproducible artifacts.
+The dataset contains groupId, artifactId, version, and the URL of all 8,292 unreproducible released modules on Maven Central.
+It also contains the reference and rebuild artifact pair of all 12,803 unreproducible artifacts.
 It is available at https://github.com/chains-project/reproducible-central/.
@@
 4.2. Methodology
-We have a dataset of 12,283 unreproducible artifacts and for each of them we have the reference artifact from Maven Central and its rebuilt version from our own rebuild.
+We have a dataset of 12,803 unreproducible artifacts and for each of them we have the reference artifact from Maven Central and its rebuilt version from our own rebuild.
 We run diffoscope 285 on each pair of reference and rebuild artifacts to get the content diff.
 diffoscope is a state-of-the-art tool for computing the difference between various types of files, archives, and directories  [^12] https://diffoscope.org/.
-This gives us a total of 12,283 diffoscope files.
-The median size of diffoscope files is 102 KB, the maximum is 3.40 GB, and the minimum is 0.38 KB.
+This gives us a total of 12,803 diffoscope files.
 Given the scale of the dataset, we first analyze a random sample of the diffoscope files to understand the reasons of unreproducibility.
-There are two attributes in the diffoscope file that are of interest to us - source1 and unified_diff, both of them indicating a reproducibility problem.
+There are two attributes in the diffoscope file that are of interest to us — source1 and unified_diff, both of them indicating a reproducibility problem.
 The source1 attribute either contains the name of the file or the command that is run to get the textual output of the file.
@@
 Another example is if source1 is javap -verbose -constants -s -l -private {}, it means that the Java class file is being compared and the reason of unreproducibility is due to differences in the bytecode.
-The unified\_diff attribute contains the diff of the file in the unified format, for textual files.
-We then write regular expressions to match these attributes to categories.
+The unified_diff attribute contains the diff of the file in the unified format, for textual files.
+Next, we notice that the changes are present in the most deeply nested unified_diff attribute which represents the textual difference in the file extracted from the archive.
+This is because Diffoscope recursively extracts files from the archive and shows the difference in the file.
+The difference in archive indicates that there is a difference in one of its file, and the textual difference in that particular file can unveil the cause of unreproducibility.
+Thus, we reduce the search space to the most deeply nested unified_diff attribute.
+This also helps us to easily read the difference in the file.
+We then write regular expressions to cluster the causes of unreproducibility based on file names and tool names reported in the source1 attribute and specific pattern in the unified_diff attribute.
 For example, if regex
 […]
-is matched, we infer that the reason of unreproducibility is due to timestamps in the file. We define regular expressions, which we match against all 12,283 diffoscope files.
-Finally, we manually go through samples that remain unmatched to the regular expressions and create expressions to map them too.
-This refines the different root causes of unreproducibility.
-Then, we report on mitigation strategies from the literature that could fix the reproducibility problem.
+is matched, we infer that the reason of unreproducibility is due to timestamps in the file and then we manually inspect the difference in the file to validate the root cause of the unreproducibility.
+The manual inspection also helps us to identify more reasons for unreproducibility.
+For example, while looking through differences in timestamps, we also observe differences about version control system information.
+This iterative process through manual analysis continuously refines the taxonomy.
+It is important to note that regular expressions are not directly mapped to the reasons we propose later, because this would have too many false positive.
+We rather rely on manual analysis to categorize differences.
+Thus, we do not count the causes of unreproducibility but build a comprehensive taxonomy about Java unreproducibility.
 4.3. Results
-Reason for Unreproducibility Root Cause of Unreproducibility Novelty Example Main Mitigation
-Build Manifests Environment - Built-By attribute Canonicalization by rebuilder
-Rebuild Process - Build-Jdk attribute Fix rebuild process
-Non-deterministic configuration Embedded branch names Fix build process
-SBOM Java Vendor Different checksum algorithms Fix rebuild process
-Custom release configuration Releasing a subset of artifacts Fix build process
-Non-deterministic information Serial Number Canonicalization by rebuilder
-Filesystem Environment - Permissions Canonicalization by rebuilder
-Custom release configuration Generated binaries are not included Fix rebuild process
-JVM Bytecode JDK Version - Ordering of constant pool entries Fix rebuild process
-Embedded data Git branch embedded Fix build process
-Build time generated code - Lambda functions Canonicalization by rebuilder
-Versioning Properties - - Number of tags embedded in Jar Canonicalization by rebuilder
-Timestamps - - Embedded in shell script in Jar files Fix build process
-Table 1. The taxonomy of unreproducibility causes based on the analysis of 12,283 unreproducible artifacts. We are the first to state the most appropriate mitigation strategy.
+Reason for Unreproducibility Root Cause of Unreproducibility Fine-grained Cause Novelty Main Mitigation
+Build Manifest Environment Built-By Canonicalization by rebuilder
+Signed JARs
+Os-Version
+Eclipse Properties
+Rebuild Process Implementation-Build-Java-Vendor Fix rebuild process
+Created-By
+Originally-Created-By
+SCM-Revision
+Build-Jdk
+Dynamic Properties SCM-Git-Branch Fix build process
+Bnd-LastModified
+Inconsistent Build Configuration Order of properties and their values Canonicalization by rebuilder
+Software Bill of Materials Java Vendor Removal of hash algorithms Fix rebuild process
+Inconsistent Build Configuration Addition of components Fix build process
+Removal of components
+Modification of components
+Dynamic Properties Timestamp Canonicalization by rebuilder
+SerialNumber
+External Metadata License Canonicalization by rebuilder
+Description of components
+ExternalReferences
+Filesystem Environment Permissions Canonicalization by rebuilder
+Ownership
+Size
+Absolute Paths
+Inconsistent Build Configuration Type Fix build process
+Files removed or added
+JVM Bytecode JDK Version Debug information Fix build process
+Optimization and de-optimization
+Refactoring
+Embedded Metadata Absolute file paths, timestamps, Java/project version, Git properties, usernames Fix build process
+Generated Code Java Compiler Canonicalization by rebuilder
+Maven Plugins Fix rebuild process
+Versioning Properties Source Repository State Number of commits Canonicalization by rebuilder
+Number of Git tags
+Branch name
+Local branch name
+Timezone of commit
+Remote URL
+Timestamps Build-time Variability Documentation Fix build process
+Shell scripts
+Executable binaries
+JVM bytecode
+Build manifest
+File metadata
+Table 2. The taxonomy of unreproducibility causes based on the analysis of 12,803 unreproducible artifacts. We are the first to study the most appropriate mitigation strategy.
 In this section, we present a comprehensive taxonomy of unreproducibility in Java, based on our large-scale, systematic protocol presented in Figure 1.
-Table 1 summarizes the taxonomy, before deep-diving into each of them.
+Table 2 summarizes the taxonomy.
 The first column lists the reason for unreproducibility.
-The second column further breaks down the reason into root causes.
-The third column indicates if the root cause is novel and has not been reported in the literature before, to our knowledge.
-The fourth column gives an example of the root cause that appears in our dataset.
-Finally, the last column lists the main mitigation strategy to fix the unreproducibility, one of the three presented in subsection 2.3.
-For example, consider the row about JVM bytecode, where unreproducibility is due to changes in the JVM bytecode between the reference and rebuild artifacts.
-One of its root cause is due to embedded data in the bytecode, this cause has not been discussed before.
-The main mitigation strategy for this problem is to fix the build process by preventing build tools to embed branch names in the JVM bytecode.
-Branch names are mutable and can change over time.
-Hence embedding them in the bytecode can lead to difference in subsequent builds.
-The other two mitigation strategy would also work, but they have cons - 1) third-party rebuilder would have to debug the build process to understand what build tool is responsible for embedding and 2) branch names have no conventional names and hence canonicalization would risk removing meaningful differences in the bytecode.
-Next, We describe each reason and its root cause in details in the following subsections.
+The second column lists the root cause that triggered the difference.
+The third column further breaks down the root cause into fine-grained causes.
+The fourth column indicates if the root cause is novel and has not been reported for Java artifacts in the literature before, including recent work by Xiong et al.
+(Xiong et al., 2022).
+Finally, the last column suggests the main strategy to fix the unreproducibility, one of the three presented in Subsection 2.3.
+For example, consider the row where ‘JVM Bytecode’ is the reason and ‘Generated Code’ is the root cause.
+We break it down into two further causes - ‘Java Compiler’ and ‘Maven Plugins’.
+A build difference due to generated JVM bytecode by Maven plugins has not been reported in the literature before.
+However, a build difference due to Java compiler has been reported by Xiong et.
+al (Xiong et al., 2022).
+For differences due to Java compiler, we can use jNorm which canonicalizes the differences. Differences due to Maven plugins can be mitigated by fixing the rebuild process.
 4.3.1. Unreproducible Build Manifests
-We consider two manifest files in our analysis - MANIFEST.MF and pom.properties.
+We consider two manifest files in our analysis — MANIFEST.MF and pom.properties.
 The MANIFEST.MF file is a mandatory metadata file in a JAR file.
@@
 Mitigation: This non-deterministic information should be stripped out from the JAR file by rebuilder before comparison.
-See section A.1, section A.4, section A.5, and section B.3 for more attributes that follow the same problem and solution.
+See section˜A.1, section˜A.4, section˜A.5, and section˜B.3 for more attributes that follow the same problem and solution.
 The next set of differences in manifests is due to differences in the script that configures the rebuild.
@@
 Mitigation: All environment specifications must be exactly specified in the the build script by rebuilder.
-Ideally, the build script file extracts information from the MANIFEST to be reproduced.
-See section A.2, section A.3, section A.6, section A.7, and section A.8 for concrete examples.
+Otherwise, the difference is not only in the attribute but it impacts other components of the release too.
+For example, the Build-Jdk attribute in the MANIFEST.MF file is set to the version of JDK used to build the JAR file and the rebuilder should use the same version of JDK to rebuild the JAR file.
+As we will see later, a different JDK versions can lead to differences in the JVM bytecode and hence the artifact becomes unreproducible.
+See section˜A.2, section˜A.3, section˜A.6, section˜A.7, and section˜A.8 for concrete examples.
 Finally, there are reasons of differences in the manifests MANIFEST.MF and pom.properties that are due to non-deterministic behavior in the build script of the builder.
 For example, build manifests pom.properties and MANIFEST.MF can differ in order of the attributes if they are not sorted.
-Mitigation: The builder should depend upon on plugins that generate sorted data  [^12] https://github.com/apache/maven-archiver/commit/763a940540eefad74f9ba73cb5eed288dc4e639d, should not embed non-deterministic data (eg branch names) in the artifact, should ignore timestamp, and stay up to date with the latest version of plugins.
-section A.9, section A.10, section A.11, section B.1, and section B.2 describe these reasons and solutions in detail.
+Mitigation: The builder should depend on plugins that generate sorted data  [^14] https://github.com/apache/maven-archiver/commit/763a940540eefad74f9ba73cb5eed288dc4e639d, should not embed non-deterministic data (e.g., branch names, timestamps) in the artifact, and stay up to date with the latest version of plugins.
+For example, only the latest version of Maven Archiver is able to maintain a deterministic file ordering.
+section˜A.9, section˜A.10, section˜A.11, section˜B.1, and section˜B.2 describe these reasons and solutions in detail.
 4.3.2. Unreproducible Software Bill of Materials
@@
 Mitigation: The rebuilder should use the same JDK vendor and version as the builder.
-See section C.1 for more details.
-The next set of changes are due to a tailored release configuration by the builder.
-For example, the builder can choose to release only a subset of the artifacts produced by the build process which can lead to less components in the released SBOM.
-Mitigation: The builder should use a conventional build process - which is to streamline the build process only using Maven commands.
+This information may be present [^16] This is not a standard practice.
+These properties could either be embedded in MANIFEST.MF or module-info.java, or documented in the README of the source repository.
+Hence, this information may not always be available and its deduction can be adhoc.
+in the Build-Jdk and Implementation-Build-Java-Vendor attribute in the MANIFEST.MF file.
+See section˜C.1 for more details.
+The next set of changes is due to a inconsistent build configuration by the builder.
+For example, the build process may use different ad-hoc scripts or rely on local caches to fetch components.
+Mitigation: The builder should use a conventional build process, such as only using standard Maven commands if only Maven projects are present in the project.
+Caches, like local Maven repository, should also be cleared because it can contain outdated components.
 This helps the rebuilder to reproduce the build process exactly without knowing all ad hoc scripting in the build.
-See section C.2 and section C.3 for concrete examples.
-Finally, the last set of changes are due to non-deterministic information in the SBOM.
+See section˜C.2 and section˜C.3 for concrete examples.
+The third set of changes is due to volatile identifiers in the SBOM.
 CycloneDX attributes like timestamp and serialNumber are non-deterministic and hence they differ in each build.
-Mitigation: There are two possible mitigations.
-First, the builder can generate the values of these attributes so that they are deterministic.
-Second, the rebuilder can strip out these attributes from the SBOM before comparing the outputs.
-See section C.5, section C.4, and section C.6 which outlines the implementation on how to either generate deterministic values or strip out the non-deterministic values.
-Reproducible SBOMs is an open problem in the field, we are the first to report about it.
+Mitigation: These attributes only act as an identifier and hence they can be canonicalized by the rebuilder before comparison.
+See section˜C.5 and section˜C.4 give concrete examples of releases falling into this case.
+The last set of changes is due to external metadata in the SBOM.
+These attributes are fetched from external sources and hence if the external source updates metadata like license, description of component, or external references, the SBOM will be different for the subsequent builds.
+Mitigation: These three attributes can be canonicalized before comparison.
+In general, data from mutable external sources should be canonicalized.
+See section˜C.6, section˜C.7, and section˜C.8 which lists releases falling into this case.
+We are the first ones to report that SBOM can be a reason of unreproducible builds.
 4.3.3. Unreproducible Filesystem
 We observe unreproducible Maven releases due to presence or absence of files, changes in the absolute file paths and their related metadata such as permissions, ownership, size and type.
-We classify them into 2 reasons below.
+We classify them into two reasons below.
 The first set of changes are due to an influence from the environment.
-For example, the permissions of the file can vary depending on the umask of the environment of rebuilder section D.2.
-The ownership and timestamp of the file can also vary depending who and when the artifact is built section D.3.
-File sizes can vary depending on the file system used to build the artifact section D.6.
-The names of the files can vary if absolute paths are used to refer to the files in artifact section D.7.
-These names can be of generated files or they could be embedded in JVM bytecode and configuration files which makes the artifact unreproducible section D.8.
-Finally, the order of files in Jar file section G.5 can vary as the order depends upon the filesystem layout  [^14] https://reproducible-builds.org/docs/archives/.
-Mitigation: The build script should either set a fixed value or strip this information from reference and rebuild artifacts before comparison.
+For example, the permissions of the file can vary depending on the umask of the environment of rebuilder section˜D.2.
+The ownership and timestamp of the file can also vary depending who and when the artifact is built section˜D.3 section˜D.4.
+File sizes can vary due to newlines as operating systems handle it differently section˜D.6.
+The names of the files can vary if absolute paths are used to refer to the files in artifact section˜D.7.
+These names can be of generated files or they could be embedded in JVM bytecode and configuration files which makes the artifact unreproducible section˜D.8.
+Finally, the order of files in Jar file section˜G.5 can vary as the order depends upon the filesystem layout  [^17] https://reproducible-builds.org/docs/archives/.
+Mitigation: The rebuilder should either set a fixed value or strip this information from the metadata of the file.
+For example, the byte buffer of zip file can be edited to set the username and group of an entry in Jar file to an empty string.
 Finally, there are some causes due to differences in the build process by the builder and rebuilder.
@@
 For the above example, the rebuilder should incorporate commands to generate the missing binaries.
-See section D.1 and section D.5 for more concrete examples.
+An automatic way to infer the correct process is to look for different Maven profiles and activate them during the rebuild process.
+See section˜D.1 and section˜D.5 for more concrete examples.
 4.3.4. Unreproducible JVM bytecode
-We observe 898 unreproducible Maven releases due to changes in the JVM bytecode when rebuilding. The reasons are as follows.
+We observe 936 unreproducible Maven releases due to changes in the JVM bytecode when rebuilding. The reasons are as follows.
 The first set of changes are due to the use of different versions of the JDK or JDK flags used in the build and rebuild process.
-This differences are due to core compilation changes such as refactoring, ordering changes in the bytecode (specifically methods, fields, static initializers, entries in the constant pool, and array type values in annotation section G.1), optimization flags, debug flags.
+This differences are due to core compilation changes such as refactoring, ordering changes in the bytecode (specifically methods, fields, static initializers, entries in the constant pool, and array type values in annotation section˜G.1), optimization flags, debug flags.
 For example, java.nio.ByteBuffer.flip() returns a different type in JDK 8 and JDK 11.
@@
 [^18] This is reported in the bug tracking system of OpenJDK (Gafter, 2002) which shows that this change is introduced in JDK 9.
-See section E.1, section E.2, and section E.3 for more examples.
+See section˜E.1, section˜E.2, and section˜E.3 for more examples.
 Mitigation: The rebuilder should use the same JDK version and flags as the builder.
@@
 This information changes with each rebuilder and hence the bytecode will be different making the artifact unreproducible.
-See section E.4 for example of a release where it happens.
+See section˜E.4 for example of a release where it happens.
 Finally, the last set of changes are due to build time code generation or modification.
@@
 Bytecode canonicalization (Sharma et al., 2024) is a technique to strip out non-deterministic information from the bytecode.
-Differences caused by Maven plugins can be mitigated by running the same plugins as in the original build process.
-Maven plugins can be configured to run only for certain build profiles.
-The rebuilder should document the build profiles used by the builder to reproduce the build process.
-See section E.5 for more examples.
+Maven can be configured to have multiple build profiles and this changes the artifacts produced.
+The rebuilder can infer the build profiles required to reproduce the build process by looking at the manifest file (pom.xml) of the artifact and choose the appropriate one which is designed for deployment.
+See section˜E.5 for more examples.
 4.3.5. Unreproducible Versioning Properties
 Git properties are embedded in Jar files in git.json or git.properties files.
-Maven plugin git-commit-id-maven-plugin  [^16] https://github.com/git-commit-id/git-commit-id-maven-plugin is commonly used to do that.
+Maven plugin git-commit-id-maven-plugin  [^19] https://github.com/git-commit-id/git-commit-id-maven-plugin is commonly used to do so.
 These properties include timezone of commit, remote URL, number of git tags in repository, name of build host, time of build, name of builder, email of builder, branch name, number of commits, local branch name, and total commits in the repository.
-All of them sometimes diverge.
+All of them may diverge between the builder and rebuilder.
 For example, value of git.tags vary for Maven release org.apache.drill:drill-opentsdb-storage:1.21.0 due to additional number of tags in the repository during rebuild.
@@
 However, we observe that there are 10 other ways timestamps can be embedded in the JAR file.
-In our dataset, we observe that: 1) they can be embedded in the properties file section F.1, 2) generated documentation section F.2, 3) shell scripts section F.3, 4) executable binaries section F.4, 5) software bill of materials section F.5, 6) JVM bytecode section F.6, 7) file metadata section F.7, 8) MANIFEST.MF section F.8, and 9) NOTICE section F.9.
-10) servlets created by Jasper JSP compiler section F.10.
+In our dataset, we observe that: 1) they can be embedded in the properties file section˜F.1, 2) generated documentation section˜F.2, 3) shell scripts section˜F.3, 4) executable binaries section˜F.4, 5) software bill of materials section˜F.5, 6) JVM bytecode section˜F.6, 7) file metadata section˜F.7, 8) MANIFEST.MF section˜F.8, and 9) NOTICE section˜F.9.
+10) servlets created by Jasper JSP compiler section˜F.10.
 Mitigation: The solution proposed by Maven is partial, more engineering is needed to support project.build.outputTimestamp in all Maven plugins that create the above mentioned files.
 4.3.7. Miscellaneous Unreproducibility Reasons
-We report reasons of reproducibility that are too specific to be categorized in the above sections.
-The first reason is due to the use of different build tools.
-For example, the Maven build tool is available as standalone tool and it also integrated into the Eclipse IDE.
-They embed information differently.
-m2e.projectLocation and m2e.projectName is embedded in pom.properties of org.spdx:spdx-maven-plugin:0.7.0 and org.owasp.antisamy:antisamy:1.7.3 section B.3.
-Even if the build tool is same, there are differences in rebuild due to usage of versions ranges of a plugin.
-For example, org.apache.aries:org.apache.aries.jax.rs.whiteboard:2.0.2 uses [1.2.5,) as the version range for flatten-maven-plugin in the POM file.
-This causes difference in the ouput of the plugin.
-Mitigation: The rebuilder should use the same build tool and its precise version as the builder.
-Moreover, the builder should avoid using version ranges for plugins in the POM file.
-The next reason is due to a difference in the operating system.
-The reference version of org.apache.helix:helix-front:1.0.3 produces a Mach-O object file format (Mac) which is different from the ELF binary format produced by the rebuilt version (Linux).
-See https://github.com/chains-project/reproducible-central/issues/20#issuecomment-2583944777 for detailed diff.
-Mitigation: The builder should ensure that 1) the OS is specified in the build specification or 2) that the binary format is independent of the operating system used to build the artifact.
-Usage of Universally Unique Identifier (UUID) is another reason of unreproducibility.
-We observe that Maven release org.apache.karaf.examples:karaf-docker-example-static-dist:4.4.0 uses UUID to generate the configuration files.
-Mitigation: The builder should use seed values to generate the UUID so that the UUID can be reproduced.
-Maven releases org.apache.helix:helix-front:1.0.3, io.wcm:io.wcm.caconfig.editor:1.9.0, and org.apache.isis.viewer:isis-viewer-restfulobjects-viewer:2.0.0-M7 embed bundled JavaScript files in the Jar file.
-These files are subject to changes in variable names and new lines.
-Mitigation: The changes in variable names can be fixed by pinning the version of the bundler.
-For example, Goswami et al.
-(Goswami et al., 2020) show that UglifyJS can produce output with different variable names for the same input if the version of UglifyJS is different.
-The difference in new lines can be fixed by the rebuilder by canonicalizing them.
-There is one instance where microbenchmarking results are embedded in the Jar file section G.4.
-Mitigation: The builder should not embed microbenchmarking results in the Jar file as they are solely for the purpose of improving performance (Costa et al., 2021) of the application and not needed for the end user.
-Finally, there are some indices files created by tools like Jandex  [^19] https://smallrye.io/jandex/jandex/3.2.1/index.html - a space efficient Java class file indexer, EFXToolkit  [^20] https://github.com/OP-TED/efx-toolkit-java - a toolkit for expressing eForms business rules, and SpotlessFormatter  [^21] https://github.com/diffplug/spotless - a code formatter for Java.
-Mitigation: The builder should not embed these index files in the artifact as they are not required for the end user and they serve only as cache files for plugins that use them.
+We report reasons of reproducibility that are specific to a particular released module. They do not fit into the categories above, so we report them here with suggested mitigations.
+Version Ranges Only one release org.apache.aries:org.apache.aries.jax.rs.whiteboard:2.0.2 is observed to use version range for flattened-maven-plugin which affects the generation of the build manifest, pom.xml.
+It is better to pin exact versions as it contributes to reproducible builds and is also a good security practice (Javan Jafari et al., 2023).
+UUIDs: Non-deterministic UUID generation in configuration files prevents reproducibility.
+JavaScript Bundling: Bundled JS files can have different variable names and line endings across builds. This has also been reported in the literature (Goswami et al., 2020).
+Microbenchmarks: Performance testing results should not be embedded in release artifacts.
+Index Files: Cache files from tools like Jandex, EFX Toolkit, and Spotless Formatter are not needed for end users.
 4.3.8. Mysteries
-We observe changes in NOTICE, POM, and other configuration files that are not due to any of the above reasons and thus, we are unable to propose a mitigation strategy for them.
-Since these reasons of unreproducibility are so diverse, we discuss them one by one in appendix H.
-{mdframed}
+We observe changes in NOTICE, build manifests, and other configuration files that are not due to any of the above reasons and thus, we do not know what caused them and cannot propose a mitigation strategy for them.
+Hence, we classify them under this section.
+We discuss the changes one by one in appendix˜H.
+4.4. Key Novelty of the Taxonomy
+Table 2 emphasizes the fine-grained causes for build unreproducibility that have never been studied in the literature before.
+Each ✓ indicates a novel finding in our work while empty cells represent causes that have already been reported in the literature, in particular by Xiong et al.
+(Xiong et al., 2022).
+Most of the categories that we have identified are not present in the related work  (Xiong et al., 2022).
+The high-level reasons for unreproducibility related to SBOM and Filesystem are not mentioned at all in the related work.
+Other categories are present in the work of Xiong et al., but we identify finer subcategories.
+For example, we report nine more subcategories for build manifest.
+Finally, a fundamental difference between our work and Xiong et al.
+(Xiong et al., 2022) is that we perform a comparison between reference and rebuild versions of the same artifact while they build twice and compare.
+We have described these fundamentally different approaches in section˜8.2.
 Answer to RQ1: “What are the causes of unreproducible builds in Java?”
 We identify 6 main reasons of unreproducibility in Java and their mitigation: 1) build manifests, 2) software bill of materials, 3) filesystem, 4) JVM bytecode, 5) versioning properties, and 6) timestamps.
-They represent the key challenges that researchers will address.
-For practitioners, they provide clear guidelines to diagnose and mitigate their unreproducible builds in Java.
+We also report novelty of our findings in Table 2 with respect to the related work.
 5. RQ2: Unreproducibility Mitigation with Artifact Canonicalization
@@
 For example, it canonicalizes file order, modification timestamp, compression algorithm, encoding, and other miscellaneous attributes from a ZIP archive.
-Chains-Rebuild  [^23] https://github.com/chains-project/chains-rebuild/commit/6dd67d5c7ac4db112f3419b5132d8f80a22cbe65 is our own fork of OSS-Rebuild, with crucial improvements dedicated to Maven reproducibility, which is unfortunately missing in the upstream project.
-We add support for canonicalizing build manifests and embedded versioning properties in JAR files as suggested in taxonomy (Table 1).
-For example, the Built-By attribute in the MANIFEST.MF file is sometimes updated to the username of the user who built the JAR file, a spurious change which is removed by Chains-Rebuild.
+Chains-Rebuild  [^23] https://github.com/chains-project/chains-rebuild/commit/6dd67d5c7ac4db112f3419b5132d8f80a22cbe65 is our own fork of OSS-Rebuild, with crucial improvements dedicated to Maven reproducibility.
+We add support for canonicalizing build manifests and embedded versioning properties in JAR files as suggested in our taxonomy (Table 2).
+For example, the Built-By attribute in the MANIFEST.MF file is sometimes updated to the username of the user who built the JAR file, a spurious change which is removed by Chains-Rebuild by directly editing the byte buffer of the JAR file.
 For versioning properties, we remove the git.properties files embedded in the JAR file.
 Finally, we remove the pom.properties file.
-We run OSS-Rebuild and Chains-Rebuild on the 12,283 reference and rebuild artifact pairs in order to analyze how effectively they make the build reproducible, that is when artifacts are bit-by-bit identical after canonicalization as shown below.
-$canonicalize(ReferenceArtifact)=canonicalize(RebuildArtifact)$
+These additional canonicalization features have been proposed to the OSS-Rebuild developers through 6 Pull Requests (#339, #360, #363, #377, #414, and #413) .
+We run OSS-Rebuild and Chains-Rebuild on the 12,803 reference and rebuild artifact pairs in order to analyze how effectively they make the build reproducible by canonicalization.
 5.3. Results
@@
 Canonicalization Canonicalization
-OSS-Rebuild (4ef4c01) 1165 (9.48%) 11118 (90.52%)
-Chains-Rebuild (6dd67d5) 3036 (24.72%) 9247 (75.28%)
-Table 2. Effectiveness of OSS-Rebuild and Chains-Rebuild on mitigating 12,283 unreproducible artifacts via canonicalization.
-Table 2 shows the results of OSS-Rebuild and Chains-Rebuild on 12,283 unreproducible artifacts.
+OSS-Rebuild (4ef4c01) 1,205 (9.41%) 11,598 (90.59%)
+Chains-Rebuild (6dd67d5) 3,405 (26.60%) 9,398 (73.40%)
+Table 3. Effectiveness of OSS-Rebuild and Chains-Rebuild on mitigating 12,803 unreproducible artifacts via canonicalization.
+Table 3 shows the results of OSS-Rebuild and Chains-Rebuild on 12,803 unreproducible artifacts.
 The first column is the name of the tool and its commit hash we use to canonicalize the artifacts.
@@
 We leave this as future work as it is purely an engineering task.
-{mdframed}
 Answer to RQ2: “To what extent does artifact canonicalization using OSS-Rebuild make a Maven release reproducible?”
-Our publicly available prototype Chains-Rebuild can canonicalize (3036) 24.71% of the artifacts.
-Our study is the first large-scale experiment on artifact canonicalization, demonstrating that it is a promising mitigation of build unreproducibility in Maven.
+Our publicly available prototype Chains-Rebuild can canonicalize (3,405) 26.60% of the artifacts from our dataset.
+Our study is the first large-scale experiment on artifact canonicalization, demonstrating that it is a promising mitigation of build unreproducibility..
 6. RQ3: Unreproducibility Mitigation with Bytecode Canonicalization
@@
 Meanwhile, in Java 5 and lower versions, it is implemented as subtraction of two integers.
-Hence, if the program is compiled with Java 5 and 6, the two bytecode representations differ even though the program is semantically the same - performing subtraction of two integers.
+Hence, if the program is compiled with Java 5 and 6, the two bytecode representations differ even though the program is semantically the same — performing subtraction of two integers.
 Canonicalization removes these differences by converting the bytecode to an intermediate form which is free from the implementation details that vary across different versions of the compiler.
@@
 Note that they refer to “canonicalization” as “normalization” in their tool.
-We run jNorm version 1.0.0 on the 898 reference rebuild artifact pairs that have differences in the bytecode.
-For maximum canonicalization, we run the tool with all the command line options as listed in the README 242424jNorm tool - https://github.com/stschott/jnorm-tool/tree/cec4645c5c9b52f73c347349bf14945b0eb55c87 of the repository and tool’s help output.
+We run jNorm version 1.0.0 on the 936 reference rebuild artifact pairs that have differences in the bytecode.
+For maximum canonicalization, we run the tool with all the command line options as listed in the README 242424jNorm tool — https://github.com/stschott/jnorm-tool/tree/cec4645c5c9b52f73c347349bf14945b0eb55c87 of the repository and tool’s help output.
 As a result of canonicalizing the artifacts, jNorm produces files in the Jimple (Raja Vallée-Rai and Laurie J.
@@
 6.3. Results
-Out of 898 artifacts, jNorm’s execution leads to the following distribution: it successfully canonicalizes 267 (29.7%) artifacts which means it produces two identical artifacts; it leaves 478 (53.2%) artifacts with a diff after canonicalization; and there are 153 (17.1%) crashes, which are ignored.
+Out of 936 artifacts, jNorm’s execution leads to the following distribution: it successfully canonicalizes 275 (29.7%) artifacts which means it produces two identical artifacts; it leaves 508 (53.2%) artifacts with a diff after canonicalization; and there are 153 (17.1%) crashes, which are ignored.
 Our results are in contrast with the results of Schott et al.
@@
 In our case, the reference and rebuild artifacts are built on different machines and based on our results in Section 4, we know that the build process is different.
-Next, we study the reasons why jNorm fails to produce identical artifacts when run on 478 cases.
+Next, we study the reasons why jNorm fails to produce identical artifacts when run on 508 cases.
 We categorize them based on the features in the bytecode that are not properly canonicalized by jNorm.
@@
     10 +return 1;
-    Listing 2: jNorm diff for control flow differnce.
+    Listing 2: jNorm diff for control flow difference.
   • Second, Maven release io.fabric8:kubernetes-model-apiextensions:6.4.0 contains generated code where the order of if-else blocks is different across different builds.
@@
 For instance, Chains-Rebuild and jNorm can be used together to canonicalize the JVM bytecode and the JAR files of a Maven artifact.
-Since Chains-Rebuild can canonicalize 3,036 artifacts and jNorm can canonicalize 267 artifacts, we can combine the 26.89% (3,303 / 12,283).
-Out of 3,303 artifacts, 2,578 artifacts are canonicalized such that 2,263 / 7,961 (28.43%) releases become fully reproducible.
-{mdframed}
+Since Chains-Rebuild can canonicalize 3,405 artifacts and jNorm can canonicalize 275 artifacts, we can combine the 28.74% (3,680 / 12,803).
+Out of 3,680 artifacts, 2,939 artifacts are canonicalized such that 2,493/ 8,292 (30.06%) releases become fully reproducible.
 Answer to RQ3: “To what extent does bytecode canonicalization with jNorm mitigate unreproducible builds in Java?”
-jNorm is able to canonicalize 267 (29.7%) out of 898 bytecode artifacts.
-However, there are 478 (53.2%) artifacts that are still unreproducible after canonicalization which calls for improvements.
-We present 4 limitation of bytecode canonicalization related to structural change, control flow, embedded data, and optimization.
-Finally, our experiments demonstrate that combining artifact and bytecode canonicalization successfully mitigates unreproducibility in 26.89% (3,303 / 12,283) artifacts.
+jNorm is able to canonicalize 275 (29.38%) out of 936 bytecode artifacts.
+However, there are 508 (54.27%) artifacts that are still unreproducible after canonicalization.
+We present in detail four limitation of bytecode canonicalization related to structural change, control flow, embedded data, and optimization.
+Finally, our experiments demonstrate that combining artifact and bytecode canonicalization successfully mitigates unreproducibility in 28.74% (3,680 / 12,803) artifacts.
+7. Best Practices for Builders and Rebuilders
+7.1. Builders
+Builders can improve reproducibility by ensuring deterministic build inputs, such as pinning dependency versions (), clearing caches before deployment, and using standard build commands rather than ad hoc scripts.
+They should avoid embedding non-deterministic data like branch names, timestamps, usernames, or UUIDs in manifests and properties, and instead use plugins that enforce deterministic ordering.
+To support verification, builders should also document and publish the exact JDK version and vendor used for releases, and generate SBOMs in a consistent and reproducible manner.
+7.2. Rebuilders
+Rebuilders should strive to mirror builder environments as closely as possible by using the same JDK vendor and version, along with consistent settings for locale, time zone, and operating system.
+When unreproducibility persists, they can apply canonicalization tools such as OSS-Rebuild and jNorm to remove spurious differences in bytecode and artifacts.
+These tools are carefully crafted to be able to canonicalize unreproducibility issues in the bytecode and artifacts.
+The remanining diff after first canonicalization can help as a feedback to the builder to fix the unreproducibility issue or to write more canonicalization rules.
 8. Related Work
@@
 Both the Tor Browser and Tails operating system emphasize that Reproducible Builds are necessary for software geared towards privacy and security.
-The Reproducible Builds  [^30] https://reproducible-builds.org/ project works on the reproducibility of coreboot  [^31] https://tests.reproducible-builds.org/coreboot/coreboot.html - an open-source firmware for x86 and ARM systems, Debian  [^32] https://tests.reproducible-builds.org/debian/reproducible.html - a Linux distribution, FreeBSD  [^33] https://tests.reproducible-builds.org/freebsd/freebsd.html and NetBSD - a UNIX based operating systems, and OpenWrt  [^34] https://tests.reproducible-builds.org/openwrt/openwrt.html - operating system for embedded devices.
+The Reproducible Builds  [^30] https://reproducible-builds.org/ project works on the reproducibility of coreboot  [^31] https://tests.reproducible-builds.org/coreboot/coreboot.html — an open-source firmware for x86 and ARM systems, Debian  [^32] https://tests.reproducible-builds.org/debian/reproducible.html — a Linux distribution, FreeBSD  [^33] https://tests.reproducible-builds.org/freebsd/freebsd.html and NetBSD — a UNIX based operating systems, and OpenWrt  [^34] https://tests.reproducible-builds.org/openwrt/openwrt.html — operating system for embedded devices.
 In the Java ecosystem, Eclipse Temurin, a vendor for Java Developement Kit, provides reproducible builds for their JDK (Leonard, 2024).
@@
 Apart from uncovering the causes of unreproducibility, they also do a survival analysis of packages to see how long it takes to fix the unreproducibility and how long the package stays reproducible.
-They show that Arch Linux pacakges become reproducible sooner than Debian packages.
+They show that Arch Linux packages become reproducible sooner than Debian packages.
 However, a Debian package is likely to stay reproducible for longer before it becomes unreproducible again.
@@
 Shi et al.
-(Shi et al., 2022) investigate reproducibility for 4 commercial systems running in Huawei and CentOS - a production grade Linux operating system.
+(Shi et al., 2022) investigate reproducibility for four commercial systems running in Huawei and CentOS — a production grade Linux operating system.
 Pöll et al.
@@
 Hence, ours paper is the first one to investigate reproducibility of normal Java applications at scale.
-The most closely related work here is the work by Xiong et al.
-(Xiong et al., 2022) that compares 59 Maven projects on Maven Central and internal registry of Huawei with the artifacts generated by the build process.
-They discover build reproducibility problems due to environment, JDK, multithreading, and other tools.
-Our work differs in the scale of the dataset (ours is 4$\times$ larger) and the taxonomy of the root causes of unreproducibility.
-Our taxonomy has more categories than the one proposed by Xiong et al.
-incl.
-new essential categories such as variations in filesystem and software bill of materials.
 Comparison between two subsequent builds
@@
 Lagnöhed (Lagnöhed, 2024) propose a tool rmake built on top of make with support for building projects twice and comparing the artifacts.
-These works differ from our work in the approach they take to verify reproducibility.
+In Java, the most closely related work here is the work by Xiong et al.
+(Xiong et al., 2022).
+They compare 59 Maven projects on Maven Central and internal registry of Huawei with the artifacts generated by the build process.
+They discover build reproducibility problems due to environment, JDK, multithreading, and other tools.
+Our work differs in the scale of the dataset (ours is 4$\times$ larger) and the taxonomy of the root causes of unreproducibility.
+Our taxonomy has more categories than the one proposed by Xiong et al.
+including new essential categories such as variations in filesystem and software bill of materials.
+All of these works differ from our work in the approach they take to verify reproducibility.
 We claim that the first approach uncovers more reasons of unreproducibility as it compares artifacts generated by build process which have maximum differences in environment.
@@
 This approach is specific for Debian packages and does not apply to Java artifacts.
-Moreover, their approach to detect unreproducibility is different from ours as mentioned in section 7.2.
+Moreover, their approach to detect unreproducibility is different from ours as mentioned in section˜8.2.
 Mukherjee et al.
@@
 Xiong et al.
-(Xiong et al., 2022) classifies fixing unreproducibility into 3 categories - 1) remediation by patching source code, third-party dependency, or build script 2) controlling which means interception of non-deterministic build instructions and returning pre-defined values for them 3) interpretation which means documenting the reason for unreproducibility.
+(Xiong et al., 2022) classifies fixing unreproducibility into three categories — 1) remediation by patching source code, third-party dependency, or build script 2) controlling which means interception of non-deterministic build instructions and returning pre-defined values for them 3) interpretation which means documenting the reason for unreproducibility.
 The authors perform these fixes automatically using a tool called JavaBEPFix which is not available to the public.
@@
 (Dietrich et al., 2024b) show that only a 25% (out of 14,156) of the binaries in Maven ecosystem are actually bit by bit identical.
-These binaries come from 4 different registries - Maven, Google Assured Open Source Software, RedHat, and Oracle build-from-source.
-They propose 4 levels of binary equivalence to compare binaries which becomes increasingly lenient as we go from level 1 to level 4.
+These binaries come from four different registries — Maven, Google Assured Open Source Software, RedHat, and Oracle build-from-source.
+They propose four levels of binary equivalence to compare binaries which becomes increasingly lenient as we go from level 1 to level 4.
 First is to compare binaries bit by bit and this is the strictest level of binary equivalence check.
@@
 Finally, strip-nondeterminism  [^38] https://salsa.debian.org/reproducible-builds/strip-nondeterminism is a tool the Debian project that strips out timestamps and file system order from archives.
-However, it only takes care of these 2 causes of unreproducibility and does not cover the other causes of unreproducibility that we have discussed in this paper.
+However, it only takes care of these two causes of unreproducibility and does not cover the other causes of unreproducibility that we have discussed in this paper.
 9. Conclusion
 In this paper, we have presented the first large-scale study of build unreproducibility in Java.
-We have identified the six main causes of unreproducibility with underlying root causes: 1) build manifest, 2) software bill of materials, 3) filesystem, 4) JVM bytecode, 5) versioning properties, and 6) timestamps.
+We have identified the six main causes of unreproducibility with underlying root causes: 1) build manifests, 2) software bill of materials, 3) filesystem, 4) JVM bytecode, 5) versioning properties, and 6) timestamps.
 Leveraging the same dataset, we have evaluated the effectiveness of canonicalization as a mitigation strategy to fix unreproducibility.
@@
 This attribute records the vendor of the JDK used to build the JAR file.
-There are 17 different open source distributions of JDK  [^39] https://sdkman.io/jdks/ - GraalVM, OpenJDK, Temurin to name a few.
+There are 17 different open source distributions of JDK  [^39] https://sdkman.io/jdks/ — GraalVM, OpenJDK, Temurin to name a few.
 Although all of them follow the same Java specification by Oracle, the implementation details could differ.
@@
 Example Artifact camel-cloud-3.13.0.jar of Maven release org.apache.camel:camel-cloud:3.13.0 is built with different versions of JDK.
-The reference version of the JAR file has the Build-Jdk attribute set to 1.8.0\_292 while the rebuild version of the JAR file has the Build-Jdk attribute set to 1.8.0\_422.
+The reference version of the JAR file has the Build-Jdk attribute set to 1.8.0_292 while the rebuild version of the JAR file has the Build-Jdk attribute set to 1.8.0_422.
 The difference in the JDK version is due to the update number (Clark and Reinhold, 2014).
@@
 Solution The attributes where this problem exists are Include-Resource, Private-Package, and Provide-Capability.
-The solution to depend upon the latest version of maven-bundle-plugin as it makes the order deterministic by sorting the values  [^43] https://github.com/apache/felix-dev/commit/d885d99a6a16660f655a4fd18e8a1a39beef0a15.
+The solution is to depend upon the latest version of maven-bundle-plugin as it makes the order deterministic by sorting the values  [^43] https://github.com/apache/felix-dev/commit/d885d99a6a16660f655a4fd18e8a1a39beef0a15.
 A.12. Addition or removal of attributes
@@
 3 #Wed Jan 29 08:18:40 UTC 2025
-Solution The timestamp should either be set to the value of SOURCE\_DATE\_EPOCH or removed from the pom.properties file.
+Solution The timestamp should either be set to the value of SOURCE_DATE_EPOCH or removed from the pom.properties file.
 B.3. Eclipse properties
-If a Maven module is built using, m2e - Maven build tool for Eclipse IDE, it adds some properties to the pom.properties file other than the ones added by standalone Maven build tool.
+If a Maven module is built using, m2e — Maven build tool for Eclipse IDE, it adds some properties to the pom.properties file other than the ones added by standalone Maven build tool.
 Example Maven release org.spdx:spdx-maven-plugin:0.7.0 has m2e.projectLocation and m2e.projectName properties in the pom.properties file.
@@
 The rebuild version documents two more Maven submodules which are net.sourceforge.pmd:pmd-cli:7.0.0 and net.sourceforge.pmd:pmd-dist:7.0.0.
-Solution The build script here  [^46] https://github.com/pmd/pmd/blob/main/.ci/build.sh#L81 shows ad-hoc configuration of the build process.
+Solution The build script here [^46] https://github.com/pmd/pmd/blob/7979570d39909195e173de05450fb1acf8b3eec5/.ci/build.sh#L81 shows ad-hoc configuration of the build process.
 This should not be done as it hinders automation when checking for reproducibility.
@@
 This has been addressed by computing serialNumber on the basis of groupId, artifactId, and version  [^52] https://github.com/CycloneDX/cyclonedx-maven-plugin/pull/425.
-C.6. System Package Data Exchange (SPDX)
+C.6. External Reference
+This attribute indicates external systems which are related to the software. For example, version control system, issue tracking system, documentation, etc.
+Example Gradle release org.jreleaser:jreleaser-mattermost-java-sdk:1.6.0 has an SBOM that only shows VCS reference for the components in the rebuilt version of the SBOM.
+Solution The external reference should be removed via canonicalization.
+C.7. Description of components
+The description attribute is a free-form text that can be used to describe the component.
+Example Gradle release org.jreleaser:jreleaser-nexus2-java-sdk:1.6.0 has an SBOM that only shows description for the components in the rebuilt version of the SBOM.
+Solution The description of the components should be removed via canonicalization.
+C.8. License
+The difference in licence appears in Software Package Data Exchange (SPDX) format.
 SPDX is a standard to document the licenses of the components used to build the software.
-It has non-deterministic attributes like created and licenseListVersion which can cause unreproducibility.
-created attribute is used to indicate the time when the SPDX file is created.
+Apart from timestamps, it has a non-deterministic attribute licenseListVersion which can cause unreproducibility.
 licenseListVersion attribute is used to indicate the version of the SPDX license list used to document the licenses of the components.
 This list is updated externally and can vary with time.
-Example We only see one instance of unreproducibility that falls under this category.
-Maven release org.apache.commons:commons-parent:60 has an artifact commons-parent-60.spdx.json where value of created and licenseListVersion attribute is different across different builds.
-Solution These attributes should be removed before comparison as they are non-deterministic.
+Example Maven release org.apache.commons:commons-parent:60 has an artifact commons-parent-60.spdx.json where value of timestamp and licenseListVersion attribute is different across different builds.
+Example Gradle release org.jreleaser:jreleaser-mattermost-java-sdk:1.6.0 has an SBOM that captures the license URLs for more components.
+Solution License attribute should be removed before comparison as they are non-deterministic.
 Appendix D Unreproducible Filesystem
@@
 If archives have different number of files, then the archive cannot be reproducible as the content of the archive is different.
-Example Maven release commons-daemon:commons-daemon:1.4.0 has an unreproducible artifact commons-daemon-1.4.0-bin-windows.zip which does not contain 3 Windows executable binary in the rebuild version of the archive.
+Example Maven release commons-daemon:commons-daemon:1.4.0 has an unreproducible artifact commons-daemon-1.4.0-bin-windows.zip which does not contain three Windows executable binary in the rebuild version of the archive.
 There is custom non-Maven command that has to be executed during the build process to generate and place these binaries under the build folder.
@@
 This corresponds to the information in the JVM bytecode that is used for debugging the Java program.
-There are three types of debug information in the JVM bytecode - SourceFile, LineNumberTable, and LocalVariableTable.
+There are three types of debug information in the JVM bytecode — SourceFile, LineNumberTable, and LocalVariableTable.
 Example In release io.dropwizard.metrics:metrics-httpclient:4.1.27, the artifact metrics-httpclient-4.1.27.jar shows an extra line in LineNumberTable which corresponds to the return JVM bytecode instruction for a void method.
@@
   • local variables in LocalVariableTable are in the scope for a more number of instructions. Example: org.apache.helix:helix-common:1.0.3.
-  • ldc\_w instructions are replaced with ldc instructions. Example: org.codehaus.mojo:jaxb2-maven-plugin:3.1.0.
+  • ldc_w instructions are replaced with ldc instructions. Example: org.codehaus.mojo:jaxb2-maven-plugin:3.1.0.
   • strings are concatenated which reduces the number of ldc instructions. Example: org.apache.sling.event:org.apache.sling.event:4.2.18.
@@
 Solution The rebuilder should sort the order independent sections of the JVM bytecode before comparison.
-Apart from order, we observe 2 types of refactoring that did not change the order.
+Apart from order, we observe two types of refactoring that did not change the order.
   • private modifier is removed from a synthetic field in an anonymous class in artifact antisamy-1.7.3.jar. We consider this as a refactoring as fields in anonymous classes are implicitly private.
@@
 Another example is org.apache.hive:kafka-handler:4.0.0-alpha-2 which uses maven-shade-plugin to rename package names from kafkaesqueesque to kafkaesque.
-Example Maven release org.apache.helix:zookeeper-api:1.0.3 has an artifact zookeeper-api-1.0.3.jar. The lambda function changes from lambda\\$parseRoutingData\\$19 to lambda\\$parseRoutingData\\$1.
+Example Maven release org.apache.helix:zookeeper-api:1.0.3 has an artifact zookeeper-api-1.0.3.jar. The lambda function changes from lambda\$parseRoutingData\$19 to lambda\$parseRoutingData\$1.
 Example Maven release org.apache.hive:hive-parser:4.0.0-alpha-2 has an artifact hive-parser-4.0.0-alpha-2-sources.jar that has sources generated by ANTLR3 Maven plugin.
@@
 This issue has been fixed by forcing the compilation of AspectJ classes  [^62] https://github.com/apache/shiro/commit/2d92460cd8b7a621be0b82725b7a5cd6952734c1.
+Example Gradle release io.opentelemetry.instrumentation:opentelemetry-instrumentation-api:2.4.0 has an artifact opentelemetry-javaagent-2.4.0.jar.diffoscope.json which records MethodParameters attribute only the in reference version of the bytecode to allow reflective access to the name of the arguments.
 Example Maven release io.fabric8:kubernetes-model
@@
 The script takes build timestamp as one of the arguments and the timestamp is different across different builds.
-Solution The shell script should respect the Maven property \${project.build.outputTimestamp} or the standard SOURCE\_DATE\_EPOCH environment variable to set the value of the timestamp in the shell script.
+Solution The shell script should respect the Maven property \${project.build.outputTimestamp} or the standard SOURCE_DATE_EPOCH environment variable to set the value of the timestamp in the shell script.
 F.4. Binaries
 Example Maven release io.github.albertus82:unexepack:0.2.1 embeds a Windows executable that embeds the timestamp.
-Solution The build tool for binary should respect SOURCE\_DATE\_EPOCH environment variable to set the value of the timestamp in the binary.
+Solution The build tool for binary should respect SOURCE_DATE_EPOCH environment variable to set the value of the timestamp in the binary.
 F.5. SBOM
-Unreproducibility due to attribute timestamp in the SBOM as explained in section C.4.
+Unreproducibility due to attribute timestamp in the SBOM as explained in section˜C.4.
 F.6. JVM bytecode
@@
 F.7. File timestamps
-Files have timestamp for creation or modification time as shown in section D.4.
+Files have timestamp for creation or modification time as shown in section˜D.4.
 F.8. MANIFEST
-Attribute Bnd-LastModified tends to be non-deterministic as documented in section A.10.
+Attribute Bnd-LastModified tends to be non-deterministic as documented in section˜A.10.
 F.9. NOTICE
@@
 G.7. Attribute values in MANIFEST
-Some values are comma separated lists whose order can differ as shown in section A.11.
+Some values are comma separated lists whose order can differ as shown in section˜A.11.
 Appendix H Mysteries
@@
 For example, "ErrordeterminingJDBCtypeforcolumn"+column+".Cause:"+e is a string concatenation expressions in artifact mybatis-3.5.16.jar.
-jNorm identifies that there are two ways to implement string concatenation (Shipilev, 2024) - using StringBuilder (JDK < 9) and using invokedynamic (JDK $\geq$ 9) instruction and canonicalizes the expression to the second one.
+jNorm identifies that there are two ways to implement string concatenation (Shipilev, 2024) — using StringBuilder (JDK < 9) and using invokedynamic (JDK $\geq$ 9) instruction and canonicalizes the expression to the second one.
 However, it fails to canonicalize the string conversion of e which is an instance of java.sql.SQLException before concatenation using invokedynamic.
@@
 jNorm is unable to handle differences in how outer class reference is handled in the bytecode.
-A synthetic field called this\\$0 is created in the bytecode to refer to the outer class from inner class.
+A synthetic field called this\$0 is created in the bytecode to refer to the outer class from inner class.
 The reference version of com.google.guava:guava:32.0.1-jre is built using custom JDK which omits this synthetic field  [^73] https://bugs.openjdk.org/browse/JDK-8271717 as it is not needed.
 We also observe that patch versions of JDK 17 handle this differently.
-They refer to the outer class reference using either this.this\\$0 or this\\$0.
+They refer to the outer class reference using either this.this\$0 or this\$0.
 The former adds another getfield instruction to the bytecode.